### v1

# FADPNet: Frequency-Aware Dual-Path Network for Face Super-Resolution

Abstract
Face super-resolution (FSR) under limited computational costs remains an open problem. Existing approaches typically treat all facial pixels equally, resulting in suboptimal allocation of computational resources and degraded FSR performance. CNN is relatively sensitive to high-frequency facial features, such as component contours and facial outlines. Meanwhile, Mamba excels at capturing low-frequency features like facial color and fine-grained texture, and does so with lower complexity than Transformers. Motivated by these observations, we propose FADPNet, a Frequency-Aware Dual-Path Network that decomposes facial features into low- and high-frequency components and processes them via dedicated branches. For low-frequency regions, we introduce a Mamba-based Low-Frequency Enhancement Block (LFEB), which combines state-space attention with squeeze-and-excitation operations to extract low-frequency global interactions and emphasize informative channels. For high-frequency regions, we design a CNN-based Deep Position-Aware Attention (DPA) module to enhance spatially-dependent structural details, complemented by a lightweight High-Frequency Refinement (HFR) module that further refines frequency-specific representations. Through the above designs, our method achieves an excellent balance between FSR quality and model efficiency, outperforming existing approaches.
Index Terms:
Face super-resolution, Mamba, Frequency feature, Dual-Path, Efficiency.
I Introduction
Face super-resolution (FSR), also known as face hallucination, focuses on reconstructing high-resolution (HR) facial images from low-resolution (LR) inputs. FSR plays a critical role in real-world scenarios where LR faces may appear in surveillance videos, historical archives, or mobile imaging. Enhancing facial image resolution is essential for improving downstream tasks such as face recognition [1], face alignment [2], and facial attribute analysis [3].
Figure 1: (a) Spectral energy distribution of features extracted by Mamba and CNN architectures, where Mamba exhibits concentrated low-frequency responses compared to CNN’s high-frequency sensitivity. Color intensity corresponds to normalized energy values. (b) Efficiency trade-offs on CelebA test set [4], demonstrating our method’s optimal balance between PSNR, Params, and inference speed.
One fundamental challenge in FSR lies in the long-tailed distribution of image frequency components [5]. High-frequency parts, such as the eyes, nose, mouth, and facial contours, account for a small portion of the facial area but require disproportionate modeling capacity due to their high variability. These regions are particularly sensitive to identity, illumination, and expression changes, making them significantly harder [6] to reconstruct than low-frequency areas such as skin, which contribute little to identity-specific features.. This imbalance results in a performance bottleneck, where failure to recover high-frequency content leads to perceptual degradation. However, most existing methods treat all pixels equally, ignoring the skewed distribution of visual complexity and importance. This uniform processing strategy leads to inefficient resource usage and suboptimal reconstruction quality. While some methods attempt to separate frequency components [7, 8], they often struggle to maintain a balanced representation across global and local facial structures.
To address these limitations, we aim to adaptively distinguish and process high- and low-frequency information. Meanwhile, we wonder about a dual-path feature fusion to help resolve the imbalance between local and global modeling that arises in previous frequency-aware methods, ultimately enabling more efficient and accurate FSR. As shown in Fig. 1(a), we present the spectral energy maps generated by CNN [9], and Mamba [10]. Interestingly, Mamba exhibits broader responses in the low-frequency regions of facial images, whereas CNN tends to be more sensitive to high-frequency components. Moreover, compared to the Transformer [11] framework, Mamba achieves global modeling through scanning mechanisms, thus keeping the computational complexity within a linear range and avoiding the expensive quadratic cost. These observations lead us to consider an alternative design: could we develop a two-branch architecture that allows Mamba and CNN to each specialize in the frequency domains they are most suited for, while integrating their outputs to better balance global and local facial structure representation?
Based on the above observations and analysis, we propose a Frequency-Aware Dual-Path Network (FADPNet), which allocates distinct submodules to specialize in different frequency domains. In particular, we introduce the Low-Frequency Enhancement Block (LFEB) and High-Frequency Enhancement Block (HFEB) to extract and refine, respectively, stable identity-preserving low-frequency features and fine-grained, structure-critical high-frequency details. The LFEB integrates an Attentive State Space Block (ASSB) and a Squeeze-and-Excitation Block (SEB) to enhance low-frequency features. At the same time, the HFEB combines a High-Frequency Refinement (HFR) module and a Depthwise Position-aware Attention (DPA) module to refine high-frequency details. These two modules integrate complementary modeling strategies—global and local—to better capture the intrinsic characteristics of facial components across different frequency levels. By decomposing and processing features based on their frequency significance, our approach not only enhances facial fidelity but also improves computational efficiency by focusing resources on perceptually important regions. As shown in Fig. 1(b), our proposed FADPNet achieves superior efficiency and reconstruction quality compared to existing FSR methods. In summary, the main contributions are as follows:
  • We propose the LFEB, which combines ASSB and SEB to enhance facial low-frequency representations by capturing global structures and emphasizing informative channel-wise responses.
  • We propose the HFEB, which combines HFR and DPA to strengthen local high-frequency facial structures and further capture long-range spatial dependencies to adaptively refine high-frequency contexts.
  • Based on the architecture designed from our frequency analysis, our FADPNet demonstrates competitive FSR ability relative to existing FSR methods across performance, model size, and inference speed.
II Related Work
II-A Face Super-Resolution
Face super-resolution [12] has made notable strides with the advancement of deep learning. Early approaches [13, 8, 14, 15] leverage explicit facial priors to guide reconstruction. For instance, FSRNet [13] incorporated facial landmarks and parsing maps to enforce structural consistency, while FAN [8] applied a face alignment network with progressive training to enhance realism. DIC [14] introduced an iterative refinement loop combining facial components and landmark predictions. To improve the representation of facial spatial features, Hu et al. [15] first utilized 3D shape priors in the network to better preserve sharp facial spatial structures.
However, prior-based methods depend heavily on accurate prior estimation, which is unreliable at extremely LR scenes and adds computational overhead. To mitigate this, recent works have adopted attention-based [16] or data-driven strategies [17]. SPARNet [16] used spatial attention to focus on key facial regions. SISN [18] decoupled attention into structural and textural streams. AD-GNN [19] combined spatial attention with graph-based message passing to model feature dependencies. LAAT [20] introduced local-aware refinement within a Transformer framework to enhance detail, while SCTANet [21] jointly modeled spatial and channel attention. SFMNet [22] employed parallel frequency-spatial branches for multi-scale feature learning. WFEN [23] used wavelet-based encoding to mitigate downsampling artifacts. However, existing FSR methods apply uniform processing across spatial regions, overlooking the varying complexity of frequency components. In contrast, our method introduces a frequency-aware dual-path design that adaptively handles high- and low-frequency features at multiple scales, leading to more robust and detail-preserving reconstruction under LR conditions.
Figure 2: Overview of our FADPNet, which adopts a U-Shape structure composed of HFEBs for high-frequency facial feature modeling and LFEBs for low-frequency facial feature enhancement.
II-B Long-Range Modeling
Modeling long-range dependencies is vital for FSR, as facial structures span distant spatial regions. Transformer-based methods [20, 22, 24] have demonstrated strong global modeling via self-attention. For example, LAAT [20] enhanced fine-grained facial modeling via self-refinement in a Transformer framework, improving detail restoration. SFMNet [22] adopted spatial and frequency branches to jointly capture global and local features across diverse receptive fields. CTCNet [25] further enhanced FSR by fusing multi-scale global representations extracted from Transformers.
However, vision Transformers [26, 25, 27] suffer from quadratic complexity, limiting their scalability. This challenge has prompted exploration of alternative long-range modeling with lower computational cost. Structured State-Space Models (SSMs)[28] have gained attention for their linear-time global modeling capability. Notably, Mamba [10] introduced a selective scan mechanism for efficient long-sequence processing while preserving strong global context, and has been successfully adapted to visual tasks [29, 30]. Building on this, MambaIR [31] extended Mamba to image restoration, alleviating issues such as local pixel forgetting and redundant channel representations. Variants like MambaLLIE [32] further demonstrate its adaptability to various low-level vision tasks.
Despite its efficiency, Mamba inherits a causal bias from its sequence-oriented design, which conflicts with the non-causal nature of super-resolution [31] and limits spatial modeling. To address this, MambaIRv2 [33] introduced direction-agnostic scanning for richer context modeling with linear scalability. Unlike existing Mamba-based methods [31, 32, 33], our approach incorporates a customized non-causal state-space backbone into the low-frequency branch, explicitly designed for face super-resolution to better preserve structural consistency and global facial context.
II-C Frequency-Aware Super-Resolution
Frequency-aware SR methods have emerged as an effective paradigm for enhancing structure, particularly in edges and textures [34, 35, 36]. By decomposing image content into distinct frequency bands, these methods enable neural networks to more precisely model the unique characteristics of each frequency component [37]. This design aligns with the well-established observation that low-frequency components mainly represent smooth and homogeneous regions, whereas high-frequency components capture fine-grained details and edge features essential for perceptual quality [38].
Explicit transformation methods decompose the image reconstruction process into frequency-specific operations. For instance, wavelet-based methods like SRCliqueNet [38] applied a discrete wavelet transform to separate LR images into sub-bands, which are then processed by specialized subnetworks to sharpen high-frequency components while preserving low-frequency structures. Similarly, FreqNet [35] operated in the discrete cosine transform domain, predicting HR frequency coefficients through learnable spectral decoders. In contrast, implicit methods achieve frequency awareness through architectural design. SRDN [39] used high-pass constrained convolutions and residual connections to emphasize high-frequency propagation, while FSN [34] adopted Octave Convolution to split features into frequency-aware streams. AFD [40] estimated feature drifting in the DCT frequency domain, adaptively enhances features. However, the above methods either concentrate solely on high-frequency enhancement [41] or fail to model the interaction between global low-frequency structures and local high-frequency details. To address this, we propose a dual-branch framework that separately models frequency-specific features, improving both reconstruction quality and computational efficiency.
III Proposed Method
III-A Overview of FADPNet
The proposed network architecture, as illustrated in Fig. 2, implements a hierarchical frequency-aware processing framework. It commences with an input image $\mathbf{I}_{LR}\in\mathbb{R}^{3\times H\times W}$ from which shallow facial features ${\mathbf{f}_{1}}\in\mathbb{R}^{C\times H\times W}$ are extracted. Our architecture adopts a three-level U-Net [42] structure, where each resolution stage includes a cascaded basic block that decomposes features into high- and low-frequency components for separate processing. The High-Frequency Enhancement Block (HFEB) restores fine-grained details using convolutions, Residual Blocks (RB), a High-Frequency Refinement (HFR) module, and a Depthwise Position-Aware Attention (DPA) module. The Low-Frequency Enhancement Block (LFEB) employs a dual-branch design: the Squeeze-and-Excitation Block (SEB) enhances local identity cues, while the Attentive State Space Block (ASSB) captures global facial structure. Outputs are fused and refined through a Feed-Forward Network (FFN). To reduce spatial misalignment caused by downsampling and upsampling, we introduce an offset-based warping module that aligns coarse structures with fine-level details, improving cross-scale consistency. Finally, features from all levels are fused and dimensionality reduction to reconstruct an HR face image $\mathbf{I}_{SR}\in\mathbb{R}^{3\times H\times W}$.
III-B Low-Frequency Enhancement Block (LFEB)
Our Low-Frequency Enhancement Block (LFEB) is designed to model low-frequency facial features, such as smooth skin regions on the forehead, cheeks, and chin, which exhibit gradual intensity variations and contain minimal identity-specific details. Specifically, given a low-frequency feature $\mathbf{F}_{L}\in\mathbb{R}^{C\times H\times W}$ extracted from the frequency separation module. First, LFEB begins with a layer normalization for stable training, followed by a dual-branch structure: the Squeeze-and-Excitation Block (SEB) [43] for adaptively calibrating feature channels to improve the focus on important features, while the Attentive State Space Block (ASSB) captures global facial structure. A learnable fusion scheme adaptively balances two feature streams, guided by scaling parameters optimized during training. The fused output is further processed by a Feed-Forward Network (FFN) [10] and combined with the block’s input via a residual connection. This design enhances identity-salient, low-frequency attributes such as coarse facial shape and alignment while preserving photorealistic appearance.
Figure 3: The architecture of Attentive State Space Block (ASSB)
Figure 4: The architecture of (a) High-Frequency Refinement (HFR) (b) Depthwise Position-aware Attention (DPA)
III-B1 Attentive State Space Block (ASSB)
As shown in Fig. 3, building upon the Mamba [11] framework, ASSB introduces non-causal global modeling through an integration of semantic perceptual sequence reorganization and a cue-guided attention mechanism, which is used to address the limitations of causal modeling in the traditional state-space. Specifically, it begins by applying a positional embedding to the input feature ${\mathbf{F}_{1}}\in\mathbb{R}^{C\times H\times W}$ to obtain a feature ${\mathbf{F}_{2}}\in\mathbb{R}^{C\times H\times W}$:
$\mathbf{F}_{2}=\mathbf{F}_{1}\cdot\sigma(\mathrm{DWConv}_{3\times 3}((\mathrm{Conv}_{\text{1}}(\mathbf{F}_{1}))),$ (1)
where $\mathrm{Conv}_{\text{1}}$ is a ${1\times 1}$ convolution ,$\mathrm{DWConv}_{3\times 3}$ denotes a ${3\times 3}$ depthwise convolution and $\sigma(\cdot)$ denotes sigmoid activation. After that, to transcend the causal constraints of standard Mamba architectures, we incorporate the learned prompts $\mathbf{P}$ into a conventional State-Space Equation (SSE), enabling richer spatial interactions beyond sequential dependency. The final prompt representation $\mathbf{P}$ consists of two components: a prompt pool $\mathbf{P}_{\text{pool}}$ and a prompt matrix $\mathbf{P}_{\text{m}}$:
${{\mathbf{P}=\mathbf{P}_{\text{m}}\cdot\mathbf{P}_{\text{pool}}}.}$ (2)
The prompt pool is parameterized via low-rank decomposition for efficiency. The shared basis $\mathbf{M}_{\text{A}}$ captures cross-block semantic commonality, while block-specific $\mathbf{M}_{\text{B}}$ enables adaptive feature combination:
$\mathbf{P}_{\text{pool}}=\mathbf{M}_{\text{A}}$ $\cdot\mathbf{M}_{\text{B}},$ (3)
$\mathbf{M}_{\text{A}}\in\mathbb{R}^{T\times r},\mathbf{M}_{\text{B}}$ $\in\mathbb{R}^{r\times d}\quad(r\ll\min\{T,d\}),$ (4)
where ${r}$ is the inner rank, ${T}$ is the number of prompts and ${d}$ is the number of hidden states in Mamba. Instance-specific prompt matrix is governed by a routing mechanism employing $\operatorname{gumbel-softmax}$ [44], which generates ${L}$ instance-specific prompts ${\mathbf{P}_{\text{m}}\in\mathbb{R}^{L\times T}}$ dynamically associated input pixels with relevant prompts from the pool:
$\mathbf{P}_{\text{m}}=\operatorname{gumbel-softmax}(\operatorname{LogSoftmax}{(\mathrm{W}_{\text{p}}\cdot\mathbf{F}_{2})),}$ (5)
where ${\mathrm{W}_{\text{p}}}$ denotes the linear layer, $\operatorname{LogSoftmax}$ refers to the logarithm of the softmax function, and gumbel-softmax is a gumbel softmax [44] operator. Then, the Semantic Guided Neighboring (SGN) unfolds this 2D feature $\mathbf{F}_{2}$ into 1D sequences for SSE. SGN-Unfold mechanism redefines the traditional spatial-to-sequential transformation by dynamically reorganizing pixels into 1D sequences based on semantic similarity, effectively mitigating the long-range decay inherent to existing models [33]. Notably, the semantic information that the mechanism relies on during this process stems from the prompt matrix $\mathbf{P}_{\text{m}}$:
$\mathrm{Index}=\operatorname{argmax}(\mathbf{P}_{\text{m}},\text{dim}=-1).$ (6)
${\mathbf{F}_{1D}=\mathrm{SNG}(\mathbf{F}_{2D},\mathrm{Index})}.$ (7)
Tokens are subsequently sorted based on their semantic indices to group semantically similar tokens into contiguous sequences. Central to ASSB operation is SSE, a modified state-space formulation that transcends the causal constraints of standard Mamba architectures. SSE enhances the output matrix ${\mathbf{C}}$ by incorporating learnable semantic prompts—compact embeddings representing pixel groups with shared semantics. This formulation injects global contextual awareness into SSE:
${{y_{i}=(\mathbf{C}+\mathbf{P})h_{i}+\mathbf{D}x_{i}}}$ (8)
Finally, an SGN-fold is applied as the inverse operation, reconstructing the semantically sorted sequence back into the original spatial layout to form the 2D feature. This reversion is guided by the previously computed reverse index, ensuring accurate spatial alignment. The restored feature is then passed through a linear projection to produce the final output feature $\mathbf{F}_{3}$. Critically, this architecture enables single-directional scanning—unlike conventional bidirectional Mamba variants requiring multiple scans—reducing computational redundancy while preserving global coherence. This efficiency breakthrough not only accelerates inference but also facilitates seamless integration with subsequent enhancement stages, where the refined low-frequency features serve as stable anchors for high-frequency detail reconstruction.
III-C High-Frequency Enhancement Block (HFEB)
In our method, HFEB specializes in high-frequency detail recovery. Specifically, for the high-frequency feature $\mathbf{F}_{H}\in\mathbb{R}^{C\times H\times W}$ obtained from frequency separation operator, we first apply a ${1\times 1}$ convolution to reduce channels to $\frac{C}{2}$ to minimizes computational costs. Then we use a ${1\times 1}$ convolution to expand channels to $C$ for follow-up operation. Subsequently, from a local enhancement perspective, features traverse a sequence of our High-Frequency Refinement (HFR), which progressively extracts hierarchical high-frequency patterns, strengthening details through iterative feature transformation. From a global modeling angle, interspersed among subsequent residual blocks, Depthwise Position-aware Attention (DPA) modules selectively enhance critical high-frequency facial attributes, while preserving facial realism through adaptive feature modulation. This synergistic operation establishes essential groundwork for achieving high-fidelity face super-resolution.
III-C1 High-Frequency Refinement (HFR)
As shown in Fig. 4 (a), our HFR module operates in a recurrent manner, to iteratively refine high-frequency components—fine-grained facial details like eyes, eyebrows, and lips—over $r$ cycles. Given an input feature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, the module first extracts multi-scale spatial features through parallel depth-wise separable convolutions with complementary receptive fields:
$\mathbf{X}_{1}=\mathrm{DWConv}_{7\times 7}(\mathbf{X}),\quad\mathbf{X}_{2}=\mathrm{DWConv}_{5\times 5}(\mathbf{X)},$ (9)
where the ${7\times 7}$ kernel captures broad contextual patterns(e.g., structural contours) and the ${5\times 5}$ kernel focuses on localized details(e.g., fine textures). These multi-scale outputs are concatenated along the channel dimension to obtain $\mathbf{X}_{\text{cat}}$:
$\mathbf{X}_{\text{cat}}=\mathrm{Concat}[\mathbf{X}_{1},\mathbf{X}_{2}].$ (10)
To integrate hierarchical spatial information. Then a ${1\times 1}$ convolution compresses the concatenated features into the original channel dimension while learning cross-scale correlations. To mitigate channel grouping-induced information isolation, a channel shuffle operation and a second ${1\times 1}$ convolution are applied to get $\mathbf{X}_{\text{3}}$:
[…]
By fusing multi-scale context and enforcing channel diversity, the HFR amplifies suppressed high-frequency signals while filtering spatial redundancies. The refined features are then expanded back to channels and fused with residual branch outputs, providing enriched high-frequency information for subsequent components in HFEB.
III-C2 Depthwise Position-aware Attention (DPA)
As shown in the Fig. 4 (b), given an input feature ${\mathbf{Y}_{1}}\in\mathbb{R}^{C\times H\times W}$, a depthwise convolution ($\mathrm{DWConv}$) is first applied to ${\mathbf{f}_{in}}$, expanding its channel dimension to $3C\times H\times W$. Following established multi-head attention frameworks, we perform channel-wise partitioning into $h$ heads, each with $\frac{C}{h}$ channels. This partitioning enables concurrent learning of distinct self-attention patterns across different feature subspaces. The partitioned feature map is then rearranged into $\mathit{Q}$, $\mathit{K}$ and $\mathit{V}$ matrices of size, defined as:
[…]
$\mathbf{pos}=\mathrm{Conv}_{7}\cdot\operatorname{GELU}(\mathrm{Conv}_{6}\cdot\mathbf{Y}_{1}),$ (13)
where $\mathrm{Conv}_{6},\mathrm{Conv}_{7}$ denote ${1\times 1}$ convolutions. However, we observe that using positional encoding alone did not yield the desired performance. To further refine the attention mechanism, we introduce a temperature generator module using a series of linear transformations to generate a temperature scale $\mathbf{temp}$, which $\mathbf{temp}$ can dynamically adjust attention scaling factors:
$\mathbf{temp}=\mathrm{W}_{2}\cdot\operatorname{GELU}(\mathrm{W}_{3}\cdot\mathbf{Y}_{1}),$ (14)
where $\mathrm{W}_{2},\mathrm{W}_{3}$ denotes the linear layer, and $\mathbf{temp}$ represents the learnable temperature parameter. Unlike fixed-temperature scaling, this adaptive mechanism provides greater flexibility in capturing long-range dependencies. In summary, the attention scores are calculated by multiplying $\mathit{Q}$ and $\mathit{K}$, and then scaled by the temperature value $\alpha$. The resulting attention-weighted $\mathit{V}$ is combined with the original positional-aware features and finally reshaped back to the original feature map size. The complete attention computation can be formulated as:
[…]
where $\sigma(\cdot)$ denotes a sigmoid activation.
III-D Offsets mechanism
As shown in Fig. 2, instead of using residual connections between input and output as in existing methods [16, 25, 23] to boost original facial information, our approach employs an offset mechanism [45] to bridge the input with the residual features of multi-scale outputs. This design stems from the observation that fixed-grid sampling in traditional convolutional or upsampling operations often struggles to capture spatially variant misalignments, especially in areas with complex geometries or motion. To tackle this, our offset mechanism adaptively predicts and compensates for these misalignments by generating pixel-wise 2D displacement vectors through a lightweight convolutional predictor subnetwork, which analyzes the input features for precise feature alignment. Specifically, for the initial input $\mathbf{f}_{1}$:
$\Delta\mathrm{~{}r=\mathbf{f}_{\text{offsets}}(\mathbf{f}_{1})\in\mathbb{R}^{2\times H\times W},\mathbf{f}_{3}=\phi(\mathbf{f}_{2},\vartriangle r),}$ (16)
where $\Delta\mathrm{r}$ is the content-related offsets matrix, $\phi(\cdot)$ denotes bilinear interpolation. Specifically, the offset-driven warping compensates for spatial misalignments caused by cascaded downsampling/upsampling operations, enhances high-frequency detail preservation through adaptively repositioning features according to local texture patterns, and promotes cross-scale feature consistency by aligning coarse-level structural information with fine-level details. Learned end-to-end without explicit supervision, this process allows the network to automatically find optimal spatial transformations that maximize feature coherence during reconstruction, which is especially crucial for keeping facial edge sharpness and structural integrity in FSR.
TABLE I: Ablation study of HFR on CelebA test [4].
Methods Params FLOPs PSNR SSIM
w/o HFR 28.16 0.8061
[…]
HFR w/ only $\mathrm{DWConv}_{5\times 5}$ 0.316M 0.689G 28.15 0.8058
HFR w/ only $\mathrm{DWConv}_{7\times 7}$ 0.390M 0.846G 28.18 0.8077
HFR 0.353M 0.802G 28.21 0.8075
TABLE II: Ablation study of DPA on Helen test set [46]. The last line denotes the strategy used in our final model.
T_gen T_para P_emb Params FLOPs PSNR SSIM
26.99 0.8034
0.162M 0.302G 26.97 0.8037
0.325M 26.93 0.8008
0.325M 0.302G 27.06 0.8064
TABLE III: Ablation study of ASSB is conducted on the Helen test set [46] by replacing it with three representative alternatives: self-attention (SA) [47], convolutional blocks (Conv), and Vision State Space Block (VSSB) [31].
[…]
SA 0.184M 0.113G 26.99 0.8035 0.4673 0.2231
CNN 0.287M 0.387G 27.01 0.8046 0.4691 0.2194
VSSB 0.269M 0.163G 27.02 0.8042 0.4697 0.2266
ASSB (Ours) 0.203M 0.125G 27.06 0.8064 0.4706 0.2167
Figure 5: Comparisons of error maps on our HFR. Redder regions indicate larger pixel-wise errors. The complete HFR module in our FADPNet leads to the lowest residual errors and shows superior ability to restore high-frequency facial details.
Figure 6: (a) Ablation studies of the effectiveness of frequency-specific modules on the CelebA [4] and Helen datasets [46]. “swap” is our model in which the low- and high-frequency blocks are swapped, resulting in a noticeable performance drop compared to our original model. (b) Feature map visualization and FSR results between our original model and swapped-frequency variant show our modules design for different frequency features produces sharp and accurate facial contours.
TABLE IV: Quantitative comparisons for $\times$8 SR on the CelebA and Helen test sets.
[…]
Bicubic 23.61 0.6779 0.1821 0.4899 22.95 0.6762 0.1745 0.4912
FSRNet [13] 27.5M 89.8ms 27.05 0.7714 0.3852 0.2127 25.45 0.7364 0.3482 0.3090
DIC [14] 22.8M 120.5ms 27.42 0.7840 0.4234 0.2129 26.15 0.7717 0.4085 0.2158
SPARNet [16] 10.6M 36.6ms 27.73 0.7949 0.4505 0.1995 26.43 0.7839 0.4262 0.2674
SISN [18] 9.8M 68ms 27.91 0.7971 0.4785 0.2005 26.64 0.7908 0.4623 0.2571
AD-GNN [19] 15.8M 107.9ms 27.82 0.7962 0.4470 0.1937 26.57 0.7786 0.4363 0.2423
Restormer-M [47] 11.7M 63.2ms 27.94 0.8027 0.4624 0.1933 26.91 0.8013 0.4595 0.2432
LAAT [20] 22.4M 35.3ms 27.91 0.7794 0.4624 0.1879 26.89 0.8005 0.4569 0.2255
ELSFace [48] 8.7M 24ms 27.96 0.7992 0.4783 0.1927 26.70 0.7910 0.4442 0.2495
SFMNet [22] 8.6M 49.2ms 27.96 0.7996 0.4644 0.1937 26.86 0.7987 0.4573 0.2322
SPADNet [49] 13.2M 52ms 27.82 0.7966 0.4589 0.1987 26.41 0.7857 0.4295 0.2645
FreqFormer [36] 9.9M 75ms 28.15 0.8047 0.4883 0.1851 26.82 0.7973 0.4678 0.2290
FADPNet (Ours) 8.6M 39.6ms 28.21 0.8075 0.4876 0.1741 27.06 0.8064 0.4706 0.2167
TABLE V: Ablation study of our attention strategy on CelebA [4] and Helen [46] test set.
Methods CelebA Helen
PSNR /  SSIM PSNR /  SSIM
LFEB w/o SEB 28.16 /  0.8067 26.99 /  0.8043
HFEB w/o DPA 28.04 /  0.8032 26.80 /  0.7981
w/o offsets 28.17 /  0.8073 27.00 /  0.8050
Ours 28.21 /  0.8075 27.06 /  0.8064
Figure 7: Visual comparisons for $\times$8 SR on the CelebA test set [4]. Our FADPNet can reconstruct clear face images.
IV Experiments
IV-A Datasets and Evaluation Metrics
We employ the CelebA [4] dataset for training and evaluate our model on three benchmarks: CelebA [4], Helen [46], and SCface [50]. Therefore, we crop the image based on its center point and resize it to ${128\times 128}$ pixels, serving as the high-resolution (HR) ground truth. Corresponding low-resolution (LR) inputs (${16\times 16}$ pixels) are generated by downsampling HR images via bicubic interpolation. The training set comprises 18,000 images from CelebA, with 1,000 and 200 images reserved for testing and validation, respectively. For cross-dataset evaluation, we test on 50 samples from Helen and the SCface dataset without fine-tuning. We assess reconstruction quality using four widely adopted metrics: PSNR and SSIM [51] for pixel-level fidelity, LPIPS [52] for perceptual similarity, and VIF [53] for texture preservation.
IV-B Implementation Details
Our FADPNet is implemented in PyTorch and trained on an $NVIDIA\quad RTX\quad 3090\quad$ GPU. We initialize the model with 32 channels and design its architecture across three distinct resolution levels. Specifically, the number of low-frequency modules is 2 at each level. Meanwhile, the number of HEFB modules is configured as follows: 2 in $\mathbb{R}^{\mathrm{C}\times\mathrm{H}\times\mathrm{W}}$ stage, 3 in $\mathbb{R}^{2\mathrm{C}\times\frac{\mathrm{H}}{2}\times\frac{\mathrm{W}}{2}}$ stage, and 4 in $\mathbb{R}^{4\mathrm{C}\times\frac{\mathrm{H}}{4}\times\frac{\mathrm{W}}{4}}$ stage. Optimization is performed using the Adam optimizer (${\beta_{1}=0.9}$ and ${\beta_{2}=0.99}$) with an initial learning rate of ${2\times 10^{-4}}$ The loss function combines reconstruction loss, and training converges within 150 epochs using a batch size of 16.
Figure 8: Visual comparisons for $\times$8 SR on the Helen test set [46]. Our FADPNet can reconstruct clear face images.
Figure 9: Visual comparisons with existing methods on downstream tasks, like face parsing.
IV-C Ablation Studies
IV-C1 Effectiveness of HFR
To further assess the impact of the HFR module in our model, we conduct an ablation study on the CelebA test set, as summarized in Table I. First, we evaluate the model without the HFR module. Removing HFR leads to a noticeable degradation in both PSNR and SSIM, indicating that the HFR module plays a crucial role in enhancing high-frequency details. We then investigate the influence of different depth-wise convolution kernel configurations. When using only the ${5\times 5}$ kernel, the model achieves a PSNR of 28.15 and SSIM of 0.8058, slightly below the performance of the configuration that employs both kernels. This suggests that the ${5\times 5}$ kernel is effective at capturing localized details but insufficient for modeling a wide range of high-frequency information. In contrast, using only the ${7\times 7}$ kernel yields a PSNR of 28.18 and SSIM of 0.8077, indicating its superiority in capturing broader contextual features. However, it still underperforms compared to the full HFR setup. Adding an extra ${3\times 3}$ kernel results in a marginal performance drop (PSNR: 28.16, SSIM: 0.8070), implying that this addition does not significantly enhance detail restoration. Moreover, removing the channel shuffle (CS) operation causes a slight decrease in both PSNR and SSIM, highlighting the importance of inter-channel interaction in feature representation. To further validate these observations, we present residual heatmaps visualizing the pixel-wise differences between the reconstructed outputs and their HR counterparts.
As shown in Fig. 5, the model without the HFR module exhibits prominent errors, especially in regions rich in high-frequency textures such as facial contours. The model with HFR but without CS shows moderate improvement, suggesting that the main structure of HFR still contributes positively even without CS. Models employing only a single kind of depth-wise convolution kernel exhibit complementary characteristics: the ${5\times 5}$ kernel better preserves localized features while missing global structures, whereas the ${7\times 7}$ kernel captures larger-scale patterns but struggles with fine textures. In contrast, the full HFR achieves the lowest errors, visually affirming its superior capacity for high-frequency detail restoration and consistent with its leading PSNR and SSIM scores.
IV-C2 Effectiveness of DPA
To comprehensively evaluate the individual and synergistic effects of components in our proposed Depthwise Position-aware Attention (DPA) module, we perform systematic ablation studies on the Helen test set [1], with detailed results presented in Table II. Our investigation focuses on three critical elements: the learnable temperature generator (T_gen), fixed temperature parameter (T_para), and positional embedding (P_emb), analyzing their impacts on both reconstruction quality and model complexity.
The initial configuration with only T_para provides a baseline level of reconstruction, but its fixed scaling restricts adaptability to spatial and contextual variations, resulting in suboptimal attention modulation for detailed facial regions. Adding P_emb to T_para slightly improves SSIM, suggesting better structural preservation, though PSNR shows a marginal decline. The P_emb component enables more precise localization of structural details, particularly in semantically critical facial regions like ocular areas, nasal contours, and lip boundaries. Our FADPNet combines T_gen with P_emb, achieving peak performance. This configuration establishes a synergistic relationship where T_gen dynamically modulates attention sharpness according to local content complexity, while P_emb ensures spatial consistency in feature aggregation. Notably, implementing T_gen without P_emb guidance degrades performance. This reveals that without spatial constraints, the self-learned scaling factors may produce inconsistent attention distributions, particularly in homogeneous facial regions where positional cues are crucial for accurate feature matching.
These systematic experiments validate the necessity of integrating both adaptive temperature scaling and explicit spatial encoding in our DPA design. The final architecture achieves superior performance through three key mechanisms: (a) Content-aware attention sharpness adaptation via T_gen, (b) Geometry-preserving feature aggregation through P_emb, and (c) Computational efficiency from depthwise implementation. The results demonstrate our module’s effectiveness in addressing the unique challenges of facial super-resolution, particularly in preserving identity-critical details while maintaining natural texture synthesis.
IV-C3 Effectiveness of ASSB
To evaluate the suitability of the Attentive State Space Block (ASSB) for low-frequency modeling in our method, we conduct ablation studies by replacing it with three representative alternatives: self-attention (SA) [47], residual block (CNN) [54], and Vision State Space Block (VSSB) [31]. As shown in Table III, our full model with ASSB achieves the best overall performance, indicating superior perceptual quality and structural fidelity.
It is worth noting that ASSB is not our original design, but a recently proposed state-space module tailored for long-range dependency modeling. In our framework, we integrate ASSB within LFEB to exploit its strength in global structure preservation. Unlike Conv, which tends to focus on local or high-frequency features, ASSB’s non-causal state-space formulation excels at capturing smooth, spatially consistent facial regions such as contours and skin tones. It’s built-in low-rank prompt decomposition also aids in suppressing high-frequency noise and promotes stable propagation across homogeneous areas. Compared to the more general-purpose VSSB, ASSB offers a lighter structure and better frequency selectivity when dealing with low-frequency signals. Despite using fewer parameters (0.203M vs. 0.269M), ASSB outperforms or matches VSSB in all metrics, demonstrating its superior efficiency-to-performance ratio when applied to low-frequency pathways. These results collectively justify our choice of ASSB as the most effective backbone for global low-frequency modeling in our frequency-aware architecture. Its integration helps maintain large-scale facial coherence and mitigates structural misalignment—challenges that conventional transformer or Conv-based designs struggle to address.
IV-C4 Study of the architecture of low- and high-frequency modules
To investigate the effectiveness of low- and high-frequency architecture, we conduct an ablation study on the CelebA and Helen datasets by interchanging the positions of the low-frequency enhancement block (LFEB) and high-frequency enhancement block (HFEB). As illustrated in Fig. 6(a), this modification leads to a consistent drop in PSNR across both datasets, highlighting the critical importance of the original frequency-aware design. The original model achieves superior reconstruction performance by appropriately allocating low- and high-frequency processing to their respective branches. In contrast, the swapped configuration disrupts this balance, resulting in suboptimal representation and degraded image quality. Fig. 6(b) visually compares the original and the swapped variants. The results clearly show that the proposed model preserves facial structures and fine details more effectively, yielding sharper contours and more realistic textures. Conversely, the swapped version produces noticeably blurrier outputs, especially around facial components like eyes and mouth, further validating the necessity of maintaining the original design of frequency-specific modules.
IV-C5 Study of local-global attention strategy
As summarized in Table V, our goal is to examine the individual contributions of local and global modeling components in both high- and low-frequency enhancement branches. First, we remove the squeeze-and-excitation block from LFEB. The performance drops on CelebA and Helen. This suggests that incorporating squeeze-and-excitation operations to extract low-frequency global interactions and emphasize informative channels is beneficial, even in smooth regions. These global interactions help preserve structurally important facial components that might otherwise be weakened in low-frequency representations. Next, we remove the DPA module from HFEB. The result shows a more noticeable degradation, with performance dropping to 28.04 / 0.8032 on CelebA and 26.80 / 0.7981 on Helen. This highlights the crucial role of DPA in capturing localized textures and reducing artifacts in high-frequency content. Finally, removing the offset mechanism causes a moderate decline in performance, indicating that it contributes to cross-scale feature alignment.
Figure 10: Visual comparison on SCface [50] of real surveillance scenarios for $\times$8 SR. Our FADPNet can reconstruct clear face components.
Figure 11: (a) PSNR, FLOPs, and speed tradeoffs on the CelebA test set [4]. (b) PSNR, FLOPs, and Params tradeoffs on the Helen test set [46], showing the balance of our method.
TABLE VI: Quantitative comparison on face average similarity with existing methods on SCface [50] of real surveillance scenarios.
Methods Average Similarity
Case 1 Case 2 Case 3 Case 4 Case 5
RCAN [54] 0.5481 0.5776 0.4441 0.7139 0.5754
FSRNet [13] 0.4930 0.5212 0.5487 0.6544 0.5276
FACN [55] 0.3670 0.3309 0.3326 0.5612 0.4373
SPARNet [16] 0.6382 0.6412 0.6455 0.6670 0.6259
SISN [18] 0.6182 0.6471 0.6354 0.7011 0.6469
AD-GNN [19] 0.6151 0.6664 0.6432 0.7213 0.6034
LAAT [20] 0.6083 0.7394 0.7282 0.7157 0.5965
SFMNet [22] 0.6240 0.6048 0.6330 0.7158 0.5742
Ours 0.6782 0.7412 0.7592 0.7714 0.6623
IV-D Comparison with Other Methods
In this section, we conduct a comprehensive comparison of our proposed method against state-of-the-art (SOTA) techniques in FSR, encompassing general image SR methods like RCAN [54], Restormer-M [47] and FreqFormer [36], specialized FSR approaches such as FSRNet [13], DIC [14], SPARNet [16], AD-GNN [19], SISN [18], LAAT [20], SFMNet [22], ELSFace [48], and SPADNet [49]. All methods are trained on the CelebA [4] dataset under identical settings—including preprocessing, loss functions, and optimizer configurations—to eliminate implementation biases and ensure direct attribution of performance differences to architectural innovations.
IV-D1 Comparison on the CelebA dataset and the Helen dataset
We present the quantitative results for the CelebA and Helen test datasets in Table IV. The best and second-best results are highlighted in bold and underlined, respectively. These results demonstrate the strength of our architecture in balancing structural accuracy and perceptual quality. To further evaluate the generalization ability of our method, we directly apply the model trained on CelebA to the Helen test set. The quantitative results for $\times$8 FSR are also presented in Table IV, showing that our method comprehensively outperforms other approaches. Despite having only 8.6M parameters (similar to SFMNet [22]), ur FADPNet achieves 39.6ms inference speed, which is 1.8× faster than FreqFormer [36] (75ms) and comparable to lightweight methods like SPARNet [16] (36.6ms). Compared to ELSFace [48], despite the similar parameter counts (8.6M vs. 8.7M), FADPNet improves PSNR by 0.25dB on CelebA and 0.36dB on Helen. As summarized in Fig. 7 and Fig. 8, while competing methods often produce blurred facial contours and distorted features (e.g., ill-defined eyes), our reconstructions successfully preserve sharp edges, facial symmetry, and fine-grained details (e.g., eyeball), closely aligning with the ground-truth HR images. As shown in Fig. 9, we also use the face parsing model [56] to analyze the FSR results of different models on the Helen dataset. The visual results clearly demonstrate that our method produces the most complete and accurate facial structure, preserving better face details than competing models. In summary, while existing methods struggle with blurred shapes and significant detail loss, our model effectively preserves facial contours and fine details. This highlights not only the robustness of our approach across different datasets but also its superiority over SOTA techniques in both performance and efficiency.
IV-D2 Comparison on real-world surveillance scenarios
Restoring facial details from real-world surveillance imagery remains a formidable challenge due to inherent low resolution, uncontrolled lighting, and sensor noise—factors that are often inadequately captured in synthetic benchmarks. To address this, we conduct experiments using low-quality face images from the SCface dataset [50], which inherently represents real-world surveillance scenarios without manual downsampling. As illustrated in Fig. 10, visual comparisons reveal that existing methods often struggle to reconstruct critical facial features due to the challenges posed by noisy, low-resolution inputs, leading to blurred contours and loss of texture details. In contrast, our FADPNet leverages adaptive feature separation and dual frequency enhancement, effectively restoring sharper facial structures, finer textures, and more natural facial symmetry. Beyond visual fidelity, we assess the impact on downstream tasks such as face matching, comparing the similarity between restored faces and their corresponding high-definition reference images. Table VI presents the quantitative results, demonstrating that our approach consistently achieves higher similarity scores across multiple test cases. This confirms its practical applicability in challenging surveillance scenarios.
IV-D3 Model Complexity Analysis
Beyond reconstruction quality, we conduct a comprehensive evaluation of model efficiency in terms of parameter count and inference speed, both of which are crucial for devices with limited computing resources. At each stage, inputs are dynamically decomposed into high-frequency and low-frequency components, which are then processed by our specialized sub-networks: a CNN branch tailored for capturing high-frequency facial details, and a Mamba-based branch better suited for modeling low-frequency, long-range dependencies. This innovative design achieves an optimal balance between computational efficiency and an enhanced representational capacity for high-frequency details. By leveraging this structured decomposition, our model adeptly handles complex facial features, leading to improved performance in FSR. As shown in Fig. 11(a) and Fig. 11(b), our method demonstrates an excellent trade-off between reconstruction accuracy and computational cost. Specifically, Fig. 11(a) illustrates that our approach surpass existing methods over 0.2dB PSNR with lower FLOPs and faster inference speed on the CelebA test set, highlighting its practical efficiency. Similarly, Fig. 11(b) further shows that on the Helen test set, our method maintains superior PSNR while keeping the parameter count relatively low, outperforming most existing methods in both accuracy and model compactness. Overall, our approach achieves a well-balanced trade-off between model size, inference speed, and reconstruction quality, making it effective.
V Limitations and Future Works
Despite the dual-path design that independently captures low- and high-frequency information, the current fusion strategy is relatively simplistic and may fail to fully exploit the complementary nature of these frequency components, resulting in suboptimal recovery of fine-grained facial details. Moreover, the model demonstrates a performance bias toward frontal faces, showing reduced robustness when handling side views or large pose variations, due to limited diversity in training views. In future work, we plan to investigate more effective fusion mechanisms between the high- and low-frequency branches to enhance inter-frequency interaction and detail restoration. Additionally, we will consider incorporating pose-aware augmentation or multi-view training strategies to improve generalization to non-frontal faces, further improving the model’s robustness in practical applications.
VI Conclusions
In this paper, we propose FADPNet, a frequency-aware dual-path network that separates and processes low- and high-frequency components for the frequency imbalance problem in face super-resolution. The Low-Frequency Enhancement Block (LFEB), comprising ASSB and SEB, models global context and reinforces identity-consistent features. In parallel, the High-Frequency Enhancement Block (HFEB), built with DPA and HFR, refines local structures through adaptive attention mechanisms. This frequency-specific design enables more effective resource allocation and promotes both structural coherence and fine-detail recovery. Extensive experiments validate the superiority of our method in the reconstruction of face image quality and computational efficiency, underscoring the effectiveness of our proposed strategy.

### v2

@@
 Abstract
-Face super-resolution (FSR) under limited computational costs remains an open problem.
-Existing approaches typically treat all facial pixels equally, resulting in suboptimal allocation of computational resources and degraded FSR performance.
-CNN is relatively sensitive to high-frequency facial features, such as component contours and facial outlines.
-Meanwhile, Mamba excels at capturing low-frequency features like facial color and fine-grained texture, and does so with lower complexity than Transformers.
-Motivated by these observations, we propose FADPNet, a Frequency-Aware Dual-Path Network that decomposes facial features into low- and high-frequency components and processes them via dedicated branches.
-For low-frequency regions, we introduce a Mamba-based Low-Frequency Enhancement Block (LFEB), which combines state-space attention with squeeze-and-excitation operations to extract low-frequency global interactions and emphasize informative channels.
-For high-frequency regions, we design a CNN-based Deep Position-Aware Attention (DPA) module to enhance spatially-dependent structural details, complemented by a lightweight High-Frequency Refinement (HFR) module that further refines frequency-specific representations.
-Through the above designs, our method achieves an excellent balance between FSR quality and model efficiency, outperforming existing approaches.
+Face super-resolution (FSR) under limited computational budgets remains challenging.
+Existing methods often treat all facial pixels equally, leading to suboptimal resource allocation and degraded performance.
+CNNs are sensitive to high-frequency facial features such as contours and outlines, while Mamba excels at capturing low-frequency attributes like facial color and texture with lower complexity than Transformers.
+Motivated by this, we propose FADPNet, a Frequency-Aware Dual-Path Network that decomposes facial features into low- and high-frequency components for dedicated processing.
+The low-frequency branch employs a Mamba-based Low-Frequency Enhancement Block (LFEB) that integrates state-space attention with squeeze-and-excitation to restore global interactions and emphasize informative channels.
+The high-frequency branch uses a CNN-based Deep Position-Aware Attention (DPA) module to refine structural details, followed by a lightweight High-Frequency Refinement (HFR) module for further frequency-specific refinement.
+These designs enable FADPNet to achieve a strong balance between FSR quality and efficiency, outperforming existing methods.
 Index Terms:
@@
 I Introduction
-Face super-resolution (FSR), also known as face hallucination, focuses on reconstructing high-resolution (HR) facial images from low-resolution (LR) inputs.
-FSR plays a critical role in real-world scenarios where LR faces may appear in surveillance videos, historical archives, or mobile imaging.
-Enhancing facial image resolution is essential for improving downstream tasks such as face recognition [1], face alignment [2], and facial attribute analysis [3].
-Figure 1: (a) Spectral energy distribution of features extracted by Mamba and CNN architectures, where Mamba exhibits concentrated low-frequency responses compared to CNN’s high-frequency sensitivity.
-Color intensity corresponds to normalized energy values.
+Face super-resolution (FSR) focuses on reconstructing high-resolution (HR) facial images from low-resolution (LR) inputs, which plays a critical role in real-world scenarios where LR faces may appear in surveillance videos, historical archives, or mobile imaging.
+Enhancing facial image resolution is essential for improving downstream tasks such as face recognition [1] and face alignment [2].
+Figure 1: (a) We divide the Fourier spectrum into low-, mid-, and high-frequency bands, and the corresponding energy ratios are averaged over the CelebA test set [3].
+Results show that Mamba preserves a noticeably larger proportion of low-frequency energy, whereas CNN retains relatively stronger high-frequency responses.
 (b) Efficiency trade-offs on CelebA test set [3], demonstrating our method’s optimal balance between PSNR, Params, and inference speed.
-One fundamental challenge in FSR lies in the long-tailed distribution of image frequency components [5].
-High-frequency parts, such as the eyes, nose, mouth, and facial contours, account for a small portion of the facial area but require disproportionate modeling capacity due to their high variability.
-These regions are particularly sensitive to identity, illumination, and expression changes, making them significantly harder [6] to reconstruct than low-frequency areas such as skin, which contribute little to identity-specific features..
-This imbalance results in a performance bottleneck, where failure to recover high-frequency content leads to perceptual degradation.
-However, most existing methods treat all pixels equally, ignoring the skewed distribution of visual complexity and importance.
-This uniform processing strategy leads to inefficient resource usage and suboptimal reconstruction quality.
-While some methods attempt to separate frequency components [7, 8], they often struggle to maintain a balanced representation across global and local facial structures.
+One core challenge in FSR is the long-tailed distribution of image frequency components [4].
+High-frequency regions, such as the eyes, nose, mouth, and facial contours, occupy only a small portion of the face yet require disproportionate modeling capacity due to their variability.
+They are highly sensitive to changes in identity and expression, making them more difficult [5] to reconstruct than low-frequency areas like skin, which contribute little to identity-specific features.
+This imbalance creates a bottleneck—failure to recover high-frequency content leads to perceptual degradation.
+However, most existing methods treat all pixels equally, ignoring the skewed distribution of visual complexity, which results in inefficient resource use and suboptimal reconstruction.
+Although some works attempt frequency separation [2], they often fail to balance global and local facial structures.
 To address these limitations, we aim to adaptively distinguish and process high- and low-frequency information.
-Meanwhile, we wonder about a dual-path feature fusion to help resolve the imbalance between local and global modeling that arises in previous frequency-aware methods, ultimately enabling more efficient and accurate FSR.
+Meanwhile, we wonder about a dual-path feature fusion to mitigate the imbalance between local and global modeling in prior frequency-aware methods.
+Low-frequency facial components are mainly associated with smooth, slowly varying, and spatially correlated patterns, such as coarse facial geometry, contour continuity, illumination distribution, and skin-tone consistency [6].
+These characteristics favor state-space modeling, whose recurrent state propagation enables efficient long-range context aggregation and promotes globally coherent representations [7, 8, 9].
+In contrast, high-frequency facial components are concentrated around localized discontinuities and rapid spatial transitions, such as eye boundaries, mouth contours, nose edges, and fine texture details [6, 10].
+Convolutions are more effective for such content because they act on local neighborhoods and respond strongly to spatial gradients and neighborhood-level variations.
+Consistent with this interpretation, we compare the band-wise frequency characteristics of intermediate features extracted by CNN [11] and Mamba [7].
 As shown in Fig.
-1(a), we present the spectral energy maps generated by CNN [9], and Mamba [10].
-Interestingly, Mamba exhibits broader responses in the low-frequency regions of facial images, whereas CNN tends to be more sensitive to high-frequency components.
-Moreover, compared to the Transformer [11] framework, Mamba achieves global modeling through scanning mechanisms, thus keeping the computational complexity within a linear range and avoiding the expensive quadratic cost.
+1(a), Mamba preserves a noticeably larger proportion of low-frequency energy, whereas CNN retains relatively stronger high-frequency responses.
 These observations lead us to consider an alternative design: could we develop a two-branch architecture that allows Mamba and CNN to each specialize in the frequency domains they are most suited for, while integrating their outputs to better balance global and local facial structure representation?
-Based on the above observations and analysis, we propose a Frequency-Aware Dual-Path Network (FADPNet), which allocates distinct submodules to specialize in different frequency domains.
-In particular, we introduce the Low-Frequency Enhancement Block (LFEB) and High-Frequency Enhancement Block (HFEB) to extract and refine, respectively, stable identity-preserving low-frequency features and fine-grained, structure-critical high-frequency details.
-The LFEB integrates an Attentive State Space Block (ASSB) and a Squeeze-and-Excitation Block (SEB) to enhance low-frequency features.
-At the same time, the HFEB combines a High-Frequency Refinement (HFR) module and a Depthwise Position-aware Attention (DPA) module to refine high-frequency details.
-These two modules integrate complementary modeling strategies—global and local—to better capture the intrinsic characteristics of facial components across different frequency levels.
-By decomposing and processing features based on their frequency significance, our approach not only enhances facial fidelity but also improves computational efficiency by focusing resources on perceptually important regions.
+Based on the above observations and analysis, we propose the Frequency-Aware Dual-Path Network (FADPNet), which allocates distinct submodules to different frequency domains.
+The Low-Frequency Enhancement Block (LFEB) integrates an Attentive State Space Block (ASSB) and Squeeze-and-Excitation Block (SEB) to restore stable, identity-preserving low-frequency features.
+The High-Frequency Enhancement Block (HFEB) combines a High-Frequency Refinement (HFR) module and Depthwise Position-aware Attention (DPA) to refine fine-grained, structure-critical details.
+These modules leverage complementary global and local modeling to capture facial characteristics across frequency levels.
+By decomposing and processing features according to frequency importance, our approach boosts facial fidelity and improves efficiency by focusing resources on perceptually critical regions.
 As shown in Fig.
-1(b), our proposed FADPNet achieves superior efficiency and reconstruction quality compared to existing FSR methods.
-In summary, the main contributions are as follows:
+1(b), FADPNet achieves superior efficiency and reconstruction quality compared to existing FSR methods.
+The main contributions of this paper are summarized as follows:
   • We propose the LFEB, which combines ASSB and SEB to enhance facial low-frequency representations by capturing global structures and emphasizing informative channel-wise responses.
@@
 Face super-resolution [12] has made notable strides with the advancement of deep learning.
-Early approaches [13, 8, 14, 15] leverage explicit facial priors to guide reconstruction.
+Early approaches [13, 2, 14, 15] leveraged explicit facial priors to guide reconstruction.
 For instance, FSRNet [13] incorporated facial landmarks and parsing maps to enforce structural consistency, while FAN [2] applied a face alignment network with progressive training to enhance realism.
@@
 [15] first utilized 3D shape priors in the network to better preserve sharp facial spatial structures.
-However, prior-based methods depend heavily on accurate prior estimation, which is unreliable at extremely LR scenes and adds computational overhead.
-To mitigate this, recent works have adopted attention-based [16] or data-driven strategies [17].
-SPARNet [16] used spatial attention to focus on key facial regions.
-SISN [18] decoupled attention into structural and textural streams.
-AD-GNN [19] combined spatial attention with graph-based message passing to model feature dependencies.
-LAAT [20] introduced local-aware refinement within a Transformer framework to enhance detail, while SCTANet [21] jointly modeled spatial and channel attention.
-SFMNet [22] employed parallel frequency-spatial branches for multi-scale feature learning.
-WFEN [23] used wavelet-based encoding to mitigate downsampling artifacts.
-However, existing FSR methods apply uniform processing across spatial regions, overlooking the varying complexity of frequency components.
-In contrast, our method introduces a frequency-aware dual-path design that adaptively handles high- and low-frequency features at multiple scales, leading to more robust and detail-preserving reconstruction under LR conditions.
+However, prior-based methods rely heavily on accurate prior estimation, which is unreliable at extremely low resolutions and adds computational overhead.
+To address this, recent works have explored attention-based [16] or data-driven [17] strategies.
+SPARNet [16] used spatial attention to focus on key facial regions, SISN [18] decoupled attention into structural and textural streams, and AD-GNN [19] combined spatial attention with graph-based message passing.
+LAAT [20] introduced local-aware refinement within a Transformer framework, SCTANet [21] jointly modeled spatial and channel attention, SFMNet [22] employed parallel frequency-spatial branches for multi-scale learning, and WFEN [23] used wavelet-based encoding to mitigate downsampling artifacts.
+Recent diffusion-based face restoration methods [24, 25], such as OSDFace [25], further demonstrated the growing interest in recovering high-quality facial structures and details with powerful generative priors.
+Nevertheless, most existing FSR methods apply uniform processing across spatial regions, ignoring the varying complexity of frequency components.
+In contrast, we propose a frequency-aware dual-path design that adaptively handles high- and low-frequency features at multiple scales, achieving more robust and detail-preserving reconstruction under LR conditions.
 Figure 2: Overview of our FADPNet, which adopts a U-Shape structure composed of HFEBs for high-frequency facial feature modeling and LFEBs for low-frequency facial feature enhancement.
 II-B Long-Range Modeling
-Modeling long-range dependencies is vital for FSR, as facial structures span distant spatial regions.
-Transformer-based methods [20, 22, 24] have demonstrated strong global modeling via self-attention.
-For example, LAAT [20] enhanced fine-grained facial modeling via self-refinement in a Transformer framework, improving detail restoration.
-SFMNet [22] adopted spatial and frequency branches to jointly capture global and local features across diverse receptive fields.
-CTCNet [25] further enhanced FSR by fusing multi-scale global representations extracted from Transformers.
-However, vision Transformers [26, 25, 27] suffer from quadratic complexity, limiting their scalability.
-This challenge has prompted exploration of alternative long-range modeling with lower computational cost.
-Structured State-Space Models (SSMs)[28] have gained attention for their linear-time global modeling capability.
-Notably, Mamba [10] introduced a selective scan mechanism for efficient long-sequence processing while preserving strong global context, and has been successfully adapted to visual tasks [29, 30].
-Building on this, MambaIR [31] extended Mamba to image restoration, alleviating issues such as local pixel forgetting and redundant channel representations.
-Variants like MambaLLIE [32] further demonstrate its adaptability to various low-level vision tasks.
-Despite its efficiency, Mamba inherits a causal bias from its sequence-oriented design, which conflicts with the non-causal nature of super-resolution [31] and limits spatial modeling.
-To address this, MambaIRv2 [33] introduced direction-agnostic scanning for richer context modeling with linear scalability.
-Unlike existing Mamba-based methods [31, 32, 33], our approach incorporates a customized non-causal state-space backbone into the low-frequency branch, explicitly designed for face super-resolution to better preserve structural consistency and global facial context.
+Modeling long-range dependencies is essential for FSR, as facial structures span distant spatial regions.
+Transformer-based methods [20, 22] achieve strong global modeling via self-attention.
+For instance, LAAT [20] refined fine-grained facial structures within a Transformer framework, SFMNet [22] combined spatial and frequency branches to capture global and local features, and CTCNet [26] fused multi-scale global representations from Transformers.
+However, vision Transformers [26, 27] incur quadratic complexity, limiting scalability.
+This has motivated exploration of lower-cost alternatives for long-range modeling.
+Structured State-Space Models offer linear-time global modeling.
+Mamba[7] introduced a selective scan mechanism for efficient long-sequence processing with strong global context, and has been adapted to visual tasks [28].
+MambaIR [8] extended Mamba to image restoration, addressing local pixel forgetting and redundant channel representations, while variants like MambaLLIE [29] show its adaptability to other low-level vision tasks.
+Despite its efficiency, Mamba inherits a causal bias from its sequence-oriented design, conflicting with the non-causal nature of super-resolution [8] and limiting spatial modeling.
+MambaIRv2 [9] alleviated this via direction-agnostic scanning for richer context modeling with linear scalability.
+DVMSR [30] demonstrated Vision Mamba’s linear-complexity global receptive field for efficient super-resolution, while MaIR [31] further validated its locality-preserving long-range modeling essential for low-frequency statistics.
+Unlike existing Mamba-based methods [8, 29, 9], we adopt a customized non-causal state-space backbone in the low-frequency branch, explicitly designed for FSR to better preserve structural consistency and global facial context.
 II-C Frequency-Aware Super-Resolution
-Frequency-aware SR methods have emerged as an effective paradigm for enhancing structure, particularly in edges and textures [34, 35, 36].
-By decomposing image content into distinct frequency bands, these methods enable neural networks to more precisely model the unique characteristics of each frequency component [37].
-This design aligns with the well-established observation that low-frequency components mainly represent smooth and homogeneous regions, whereas high-frequency components capture fine-grained details and edge features essential for perceptual quality [38].
-Explicit transformation methods decompose the image reconstruction process into frequency-specific operations.
-For instance, wavelet-based methods like SRCliqueNet [38] applied a discrete wavelet transform to separate LR images into sub-bands, which are then processed by specialized subnetworks to sharpen high-frequency components while preserving low-frequency structures.
-Similarly, FreqNet [35] operated in the discrete cosine transform domain, predicting HR frequency coefficients through learnable spectral decoders.
-In contrast, implicit methods achieve frequency awareness through architectural design.
-SRDN [39] used high-pass constrained convolutions and residual connections to emphasize high-frequency propagation, while FSN [34] adopted Octave Convolution to split features into frequency-aware streams.
-AFD [40] estimated feature drifting in the DCT frequency domain, adaptively enhances features.
-However, the above methods either concentrate solely on high-frequency enhancement [41] or fail to model the interaction between global low-frequency structures and local high-frequency details.
-To address this, we propose a dual-branch framework that separately models frequency-specific features, improving both reconstruction quality and computational efficiency.
+Frequency-aware SR methods have proven effective for enhancing structures, particularly edges and textures [32, 33].
+By decomposing image content into distinct frequency bands, they enable networks to model the characteristics of each component more precisely [34].
+Low-frequency components typically represent smooth, homogeneous regions, while high-frequency components capture fine details and edges essential for perceptual quality [6].
+Explicit transformation methods decompose reconstruction into frequency-specific operations.
+Wavelet-based methods, such as SRCliqueNet [6], used a discrete wavelet transform to split LR images into sub-bands, processed by specialized subnetworks to sharpen high-frequency components while preserving low-frequency structures.
+FreqNet [32] operated in the discrete cosine transform domain, predicting HR frequency coefficients with learnable spectral decoders.
+In contrast, implicit methods incorporate frequency awareness into network design.
+SRDN [10] employed high-pass constrained convolutions and residual connections to emphasize high-frequency propagation, AFD [35] estimated feature drifting in the DCT domain to adaptively enhance features.
+Recent works such as LSRNA [36] further improved detail fidelity in higher-resolution diffusion-based image generation through latent-space super-resolution and region-aware enhancement.
+However, existing frequency-aware methods [37] usually improve reconstruction through frequency enhancement or decomposition, while the different modeling needs of low- and high-frequency facial information are not explicitly addressed.
+In contrast, our dual-branch framework separately captures global low-frequency structures and local high-frequency details, improving both reconstruction quality and efficiency.
 III Proposed Method
 III-A Overview of FADPNet
-The proposed network architecture, as illustrated in Fig.
-2, implements a hierarchical frequency-aware processing framework.
-It commences with an input image $\mathbf{I}_{LR}\in\mathbb{R}^{3\times H\times W}$ from which shallow facial features ${\mathbf{f}_{1}}\in\mathbb{R}^{C\times H\times W}$ are extracted.
-Our architecture adopts a three-level U-Net [42] structure, where each resolution stage includes a cascaded basic block that decomposes features into high- and low-frequency components for separate processing.
-The High-Frequency Enhancement Block (HFEB) restores fine-grained details using convolutions, Residual Blocks (RB), a High-Frequency Refinement (HFR) module, and a Depthwise Position-Aware Attention (DPA) module.
-The Low-Frequency Enhancement Block (LFEB) employs a dual-branch design: the Squeeze-and-Excitation Block (SEB) enhances local identity cues, while the Attentive State Space Block (ASSB) captures global facial structure.
+The proposed architecture, shown in Fig.
+2, follows a hierarchical frequency-aware framework.
+Given an input image $\mathbf{I}_{LR}\in\mathbb{R}^{3\times H\times W}$, shallow features ${\mathbf{f}_{1}}\in\mathbb{R}^{C\times H\times W}$ are first extracted.
+We adopt a three-level U-Net, where each stage uses a cascaded basic block to decompose features into high- and low-frequency components for separate processing.
+The High-Frequency Enhancement Block (HFEB) refines fine details via convolutions, Residual Blocks (RB), a High-Frequency Refinement (HFR) module, and Depthwise Position-Aware Attention (DPA).
+The Low-Frequency Enhancement Block (LFEB) employs a dual-branch design: a Squeeze-and-Excitation Block (SEB) for local identity cues and an Attentive State Space Block (ASSB) for global structure modeling.
 Outputs are fused and refined through a Feed-Forward Network (FFN).
-To reduce spatial misalignment caused by downsampling and upsampling, we introduce an offset-based warping module that aligns coarse structures with fine-level details, improving cross-scale consistency.
-Finally, features from all levels are fused and dimensionality reduction to reconstruct an HR face image $\mathbf{I}_{SR}\in\mathbb{R}^{3\times H\times W}$.
+To mitigate misalignment from down/up-sampling, an offset-based warping module aligns coarse structures with fine details, enhancing cross-scale consistency.
+Finally, multi-level features are fused and reduced to reconstruct the HR face image $\mathbf{I}_{SR}\in\mathbb{R}^{3\times H\times W}$.
 III-B Low-Frequency Enhancement Block (LFEB)
-Our Low-Frequency Enhancement Block (LFEB) is designed to model low-frequency facial features, such as smooth skin regions on the forehead, cheeks, and chin, which exhibit gradual intensity variations and contain minimal identity-specific details.
-Specifically, given a low-frequency feature $\mathbf{F}_{L}\in\mathbb{R}^{C\times H\times W}$ extracted from the frequency separation module.
-First, LFEB begins with a layer normalization for stable training, followed by a dual-branch structure: the Squeeze-and-Excitation Block (SEB) [43] for adaptively calibrating feature channels to improve the focus on important features, while the Attentive State Space Block (ASSB) captures global facial structure.
-A learnable fusion scheme adaptively balances two feature streams, guided by scaling parameters optimized during training.
-The fused output is further processed by a Feed-Forward Network (FFN) [10] and combined with the block’s input via a residual connection.
-This design enhances identity-salient, low-frequency attributes such as coarse facial shape and alignment while preserving photorealistic appearance.
-Figure 3: The architecture of Attentive State Space Block (ASSB)
-Figure 4: The architecture of (a) High-Frequency Refinement (HFR) (b) Depthwise Position-aware Attention (DPA)
+Our Low-Frequency Enhancement Block (LFEB) models smooth facial regions (forehead, cheeks, chin) with gradual intensity variations and limited identity-specific details.
+Given low-frequency features $\mathbf{F}_{L}\in\mathbb{R}^{C\times H\times W}$ from the frequency separation module, LFEB applies layer normalization, followed by two branches: a Squeeze-and-Excitation Block (SEB)[38] for adaptive channel calibration and an Attentive State Space Block (ASSB) for global structure modeling.
+A learnable fusion scheme with trainable scaling parameters balances the streams, after which a Feed-Forward Network (FFN)[7] refines the output and adds it to the block input via a residual connection.
+This design restores coarse facial shape and alignment while preserving photorealism.
+Figure 3: Architecture of Attentive State Space Block (ASSB).
+Figure 4: The architecture of (a) High-Frequency Refinement (HFR), (b) Depthwise Position-aware Attention (DPA).
 III-B1 Attentive State Space Block (ASSB)
@@
 $\mathbf{M}_{\text{A}}\in\mathbb{R}^{T\times r},\mathbf{M}_{\text{B}}$ $\in\mathbb{R}^{r\times d}\quad(r\ll\min\{T,d\}),$ (4)
-where ${r}$ is the inner rank, ${T}$ is the number of prompts and ${d}$ is the number of hidden states in Mamba.
-Instance-specific prompt matrix is governed by a routing mechanism employing $\operatorname{gumbel-softmax}$ [44], which generates ${L}$ instance-specific prompts ${\mathbf{P}_{\text{m}}\in\mathbb{R}^{L\times T}}$ dynamically associated input pixels with relevant prompts from the pool:
+where r{r} is the inner rank, T{T} is the number of prompts and d{d} is the number of hidden states in Mamba.
+Instance-specific prompt matrix is governed by a routing mechanism employing $\operatorname{gumbel-softmax}$ [39], which generates L{L} instance-specific prompts $\mathbf{P}_{\text{m}}\in\mathbb{R}^{L\times T}$} dynamically associated with input pixels and relevant prompts from the pool:
 $\mathbf{P}_{\text{m}}=\operatorname{gumbel-softmax}(\operatorname{LogSoftmax}{(\mathrm{W}_{\text{p}}\cdot\mathbf{F}_{2})),}$ (5)
@@
 Then, the Semantic Guided Neighboring (SGN) unfolds this 2D feature $\mathbf{F}_{2}$ into 1D sequences for SSE.
-SGN-Unfold mechanism redefines the traditional spatial-to-sequential transformation by dynamically reorganizing pixels into 1D sequences based on semantic similarity, effectively mitigating the long-range decay inherent to existing models [33].
+SGN-Unfold mechanism redefines the traditional spatial-to-sequential transformation by dynamically reorganizing pixels into 1D sequences based on semantic similarity, effectively mitigating the long-range decay inherent to existing methods [9].
 Notably, the semantic information that the mechanism relies on during this process stems from the prompt matrix $\mathbf{P}_{\text{m}}$:
@@
 This formulation injects global contextual awareness into SSE:
-${{y_{i}=(\mathbf{C}+\mathbf{P})h_{i}+\mathbf{D}x_{i}}}$ (8)
-Finally, an SGN-fold is applied as the inverse operation, reconstructing the semantically sorted sequence back into the original spatial layout to form the 2D feature.
-This reversion is guided by the previously computed reverse index, ensuring accurate spatial alignment.
-The restored feature is then passed through a linear projection to produce the final output feature $\mathbf{F}_{3}$.
-Critically, this architecture enables single-directional scanning—unlike conventional bidirectional Mamba variants requiring multiple scans—reducing computational redundancy while preserving global coherence.
-This efficiency breakthrough not only accelerates inference but also facilitates seamless integration with subsequent enhancement stages, where the refined low-frequency features serve as stable anchors for high-frequency detail reconstruction.
+$y_{i}=(\mathbf{C}+\mathbf{P})h_{i}+\mathbf{D}x_{i}$}}. (8)
+Finally, an SGN-fold inversely reconstructs the semantically sorted sequence into the original 2D layout using the reverse index for precise spatial alignment.
+The restored feature is then linearly projected to produce the final output $\mathbf{F}_{3}$.
+Unlike conventional bidirectional Mamba variants, our single-directional scanning avoids multiple passes, reducing redundancy while maintaining global coherence.
+This improves inference speed and enables smooth integration with subsequent enhancement stages, where refined low-frequency features anchor high-frequency detail reconstruction.
 III-C High-Frequency Enhancement Block (HFEB)
-In our method, HFEB specializes in high-frequency detail recovery.
-Specifically, for the high-frequency feature $\mathbf{F}_{H}\in\mathbb{R}^{C\times H\times W}$ obtained from frequency separation operator, we first apply a ${1\times 1}$ convolution to reduce channels to $\frac{C}{2}$ to minimizes computational costs.
-Then we use a ${1\times 1}$ convolution to expand channels to $C$ for follow-up operation.
-Subsequently, from a local enhancement perspective, features traverse a sequence of our High-Frequency Refinement (HFR), which progressively extracts hierarchical high-frequency patterns, strengthening details through iterative feature transformation.
-From a global modeling angle, interspersed among subsequent residual blocks, Depthwise Position-aware Attention (DPA) modules selectively enhance critical high-frequency facial attributes, while preserving facial realism through adaptive feature modulation.
-This synergistic operation establishes essential groundwork for achieving high-fidelity face super-resolution.
+In our method, the High-Frequency Enhancement Block (HFEB) targets high-frequency detail recovery.
+Given high-frequency features $\mathbf{F}_{H}\in\mathbb{R}^{C\times H\times W}$ from the frequency separation operator, we first apply a $1\times 1$} convolution to reduce channels to $\frac{C}{2}$, then expand them back to $C$ to reduce computation.
+Locally, features pass through the High-Frequency Refinement (HFR) module, which progressively extracts hierarchical high-frequency patterns and strengthens details.
+Globally, Depthwise Position-aware Attention (DPA) modules interleaved with residual blocks selectively enhance critical high-frequency attributes while preserving realism.
+This combination forms the basis for high-fidelity FSR.
 III-C1 High-Frequency Refinement (HFR)
 As shown in Fig.
-4 (a), our HFR module operates in a recurrent manner, to iteratively refine high-frequency components—fine-grained facial details like eyes, eyebrows, and lips—over $r$ cycles.
+4 (a), our HFR module operates recurrently, to iteratively refine high-frequency components—fine-grained facial details like eyes, eyebrows, and lips—over $r$ cycles.
 Given an input feature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, the module first extracts multi-scale spatial features through parallel depth-wise separable convolutions with complementary receptive fields:
@@
 To mitigate channel grouping-induced information isolation, a channel shuffle operation and a second $1\times 1$} convolution are applied to get $\mathbf{X}_{\text{3}}$:
-[…]
-[…]
+$\mathbf{X}_{\text{3}}=\mathrm{Conv}_{\text{4}}\cdot\mathit{CS}(\mathrm{Conv}_{\text{3}}\cdot\mathbf{X}_{\text{cat}}),$ (11)
+where $\mathrm{Conv}_{\text{3}}$ is a $1\times 1$} convolution for channel expansion, $\mathit{CS}(\cdot)$ denotes channel-shuffle and $\mathrm{Conv}_{\text{4}}$ is a $1\times 1$} convolution for channel reduction.
+Combined with channel shuffle, this operation dynamically permutes channel subgroups to enhance inter-channel communication.
 By fusing multi-scale context and enforcing channel diversity, the HFR amplifies suppressed high-frequency signals while filtering spatial redundancies.
@@
 The partitioned feature map is then rearranged into $\mathit{Q}$, $\mathit{K}$ and $\mathit{V}$ matrices of size, defined as:
-[…]
-[…]
+$\mathit{Q,K,V}=\mathit{RS}\left(\mathrm{DWConv}_{3\times 3}\left(\mathrm{Conv}_{\text{5}}\left(\mathbf{Y}_{1}\right)\right)\right),$} (12)
+where $\mathrm{Conv}_{\text{5}}$ is a $1\times 1$} convolution that expands channels from $C$ to $3C$, $\mathrm{DWConv}_{3\times 3}$ denotes a $3\times 3$} depthwise convolution and $\mathit{RS}$ denotes a reshape and split operator.
+To enhance positional modeling capabilities, we introduce a dynamical positional encoding mechanism.
+The input feature map is processed through a sub-network consisting of two $1\times 1\times 1$ convolutions with $\operatorname{GELU}$ activation in between:
 $\mathbf{pos}=\mathrm{Conv}_{7}\cdot\operatorname{GELU}(\mathrm{Conv}_{6}\cdot\mathbf{Y}_{1}),$ (13)
 where $\mathrm{Conv}_{6},\mathrm{Conv}_{7}$ denote $1\times 1$} convolutions.
-However, we observe that using positional encoding alone did not yield the desired performance.
-To further refine the attention mechanism, we introduce a temperature generator module using a series of linear transformations to generate a temperature scale $\mathbf{temp}$, which $\mathbf{temp}$ can dynamically adjust attention scaling factors:
+However, we observe that using Positional Embedding alone did not yield the desired performance.
+To further refine the attention mechanism, we introduce a Temperature Generator module using a series of linear transformations to generate a temperature scale $\mathbf{temp}$, which can dynamically adjust attention scaling factors:
 $\mathbf{temp}=\mathrm{W}_{2}\cdot\operatorname{GELU}(\mathrm{W}_{3}\cdot\mathbf{Y}_{1}),$ (14)
@@
 The complete attention computation can be formulated as:
-[…]
+$\operatorname{Attn}\mathit{\{Q,K,V\}}=\mathit{V}\cdot\operatorname{RELU}(\mathit{QK}^{T}\cdot\mathbf{temp})+\sigma(\mathbf{pos}),$ (15)
 where $\sigma(\cdot)$ denotes a sigmoid activation.
@@
 Learned end-to-end without explicit supervision, this process allows the network to automatically find optimal spatial transformations that maximize feature coherence during reconstruction, which is especially crucial for keeping facial edge sharpness and structural integrity in FSR.
-TABLE I: Ablation study of HFR on CelebA test [4].
-Methods Params FLOPs PSNR SSIM
-w/o HFR 28.16 0.8061
-[…]
-HFR w/ only $\mathrm{DWConv}_{5\times 5}$ 0.316M 0.689G 28.15 0.8058
-HFR w/ only $\mathrm{DWConv}_{7\times 7}$ 0.390M 0.846G 28.18 0.8077
-HFR 0.353M 0.802G 28.21 0.8075
-TABLE II: Ablation study of DPA on Helen test set [46]. The last line denotes the strategy used in our final model.
-T_gen T_para P_emb Params FLOPs PSNR SSIM
-26.99 0.8034
-0.162M 0.302G 26.97 0.8037
-0.325M 26.93 0.8008
-0.325M 0.302G 27.06 0.8064
-TABLE III: Ablation study of ASSB is conducted on the Helen test set [46] by replacing it with three representative alternatives: self-attention (SA) [47], convolutional blocks (Conv), and Vision State Space Block (VSSB) [31].
-[…]
-SA 0.184M 0.113G 26.99 0.8035 0.4673 0.2231
-CNN 0.287M 0.387G 27.01 0.8046 0.4691 0.2194
-VSSB 0.269M 0.163G 27.02 0.8042 0.4697 0.2266
-ASSB (Ours) 0.203M 0.125G 27.06 0.8064 0.4706 0.2167
-Figure 5: Comparisons of error maps on our HFR.
-Redder regions indicate larger pixel-wise errors.
-The complete HFR module in our FADPNet leads to the lowest residual errors and shows superior ability to restore high-frequency facial details.
-Figure 6: (a) Ablation studies of the effectiveness of frequency-specific modules on the CelebA [4] and Helen datasets [46].
-“swap” is our model in which the low- and high-frequency blocks are swapped, resulting in a noticeable performance drop compared to our original model.
-(b) Feature map visualization and FSR results between our original model and swapped-frequency variant show our modules design for different frequency features produces sharp and accurate facial contours.
-TABLE IV: Quantitative comparisons for $\times$8 SR on the CelebA and Helen test sets.
-[…]
-[…]
-Bicubic 23.61 0.6779 0.1821 0.4899 22.95 0.6762 0.1745 0.4912
-FSRNet [13] 27.5M 89.8ms 27.05 0.7714 0.3852 0.2127 25.45 0.7364 0.3482 0.3090
-DIC [14] 22.8M 120.5ms 27.42 0.7840 0.4234 0.2129 26.15 0.7717 0.4085 0.2158
-SPARNet [16] 10.6M 36.6ms 27.73 0.7949 0.4505 0.1995 26.43 0.7839 0.4262 0.2674
-SISN [18] 9.8M 68ms 27.91 0.7971 0.4785 0.2005 26.64 0.7908 0.4623 0.2571
-AD-GNN [19] 15.8M 107.9ms 27.82 0.7962 0.4470 0.1937 26.57 0.7786 0.4363 0.2423
-Restormer-M [47] 11.7M 63.2ms 27.94 0.8027 0.4624 0.1933 26.91 0.8013 0.4595 0.2432
-LAAT [20] 22.4M 35.3ms 27.91 0.7794 0.4624 0.1879 26.89 0.8005 0.4569 0.2255
-ELSFace [48] 8.7M 24ms 27.96 0.7992 0.4783 0.1927 26.70 0.7910 0.4442 0.2495
-SFMNet [22] 8.6M 49.2ms 27.96 0.7996 0.4644 0.1937 26.86 0.7987 0.4573 0.2322
-SPADNet [49] 13.2M 52ms 27.82 0.7966 0.4589 0.1987 26.41 0.7857 0.4295 0.2645
-FreqFormer [36] 9.9M 75ms 28.15 0.8047 0.4883 0.1851 26.82 0.7973 0.4678 0.2290
-FADPNet (Ours) 8.6M 39.6ms 28.21 0.8075 0.4876 0.1741 27.06 0.8064 0.4706 0.2167
-TABLE V: Ablation study of our attention strategy on CelebA [4] and Helen [46] test set.
-Methods CelebA Helen
-PSNR /  SSIM PSNR /  SSIM
-LFEB w/o SEB 28.16 /  0.8067 26.99 /  0.8043
-HFEB w/o DPA 28.04 /  0.8032 26.80 /  0.7981
-w/o offsets 28.17 /  0.8073 27.00 /  0.8050
-Ours 28.21 /  0.8075 27.06 /  0.8064
-Figure 7: Visual comparisons for $\times$8 SR on the CelebA test set [4]. Our FADPNet can reconstruct clear face images.
+III-E Loss Function
+Given a training pair $\left\{\boldsymbol{I}_{LR},\boldsymbol{I}_{HR}\right\}$, our model predicts the super-resolved image $\boldsymbol{I}_{SR}$ from the low-resolution input $\boldsymbol{I}_{LR}$.
+The overall training objective is defined as
+$\mathcal{L}_{total}=\mathcal{L}_{pixel}=\left\|\boldsymbol{I}_{SR}-\boldsymbol{I}_{HR}\right\|_{1},$ (17)
+where $\lambda=1$, and $\boldsymbol{I}_{HR}$ denotes the corresponding high-resolution ground truth. No additional perceptual, adversarial, or frequency-domain loss is employed in training.
 IV Experiments
@@
 We assess reconstruction quality using four widely adopted metrics: PSNR and SSIM [43] for pixel-level fidelity, LPIPS [44] for perceptual similarity, and VIF [45] for texture preservation.
+For runtime evaluation, the reported runtime is computed as the average inference time over 1,000 test images with batch size 1.
+Inference is performed in FP32 precision without warmup runs, and the reported timing only includes network forward inference, excluding preprocessing and data transfer.
 IV-B Implementation Details
@@
 Specifically, the number of low-frequency modules is 2 at each level.
-Meanwhile, the number of HEFB modules is configured as follows: 2 in $\mathbb{R}^{\mathrm{C}\times\mathrm{H}\times\mathrm{W}}$ stage, 3 in $\mathbb{R}^{2\mathrm{C}\times\frac{\mathrm{H}}{2}\times\frac{\mathrm{W}}{2}}$ stage, and 4 in $\mathbb{R}^{4\mathrm{C}\times\frac{\mathrm{H}}{4}\times\frac{\mathrm{W}}{4}}$ stage.
-Optimization is performed using the Adam optimizer (${\beta_{1}=0.9}$ and ${\beta_{2}=0.99}$) with an initial learning rate of ${2\times 10^{-4}}$ The loss function combines reconstruction loss, and training converges within 150 epochs using a batch size of 16.
-Figure 8: Visual comparisons for $\times$8 SR on the Helen test set [46]. Our FADPNet can reconstruct clear face images.
-Figure 9: Visual comparisons with existing methods on downstream tasks, like face parsing.
+Meanwhile, the number of HFEB modules is configured as follows: 2 in $\mathbb{R}^{\mathrm{C}\times\mathrm{H}\times\mathrm{W}}$ stage, 3 in $\mathbb{R}^{2\mathrm{C}\times\frac{\mathrm{H}}{2}\times\frac{\mathrm{W}}{2}}$ stage, and 4 in $\mathbb{R}^{4\mathrm{C}\times\frac{\mathrm{H}}{4}\times\frac{\mathrm{W}}{4}}$ stage.
+Optimization is performed using the Adam optimizer ($\beta_{1}=0.9$} and $\beta_{2}=0.99$}) with an initial learning rate of $2\times 10^{-4}$}.
+The model is optimized using the reconstruction loss defined in Section III-E, and training converges within 150 epochs using a batch size of 16.
+During training, standard data augmentation is applied, including random horizontal flipping and random scaling with a factor sampled from $[1.0,1.3]$.
+The low-resolution input is generated by bicubic downsampling of the HR image, followed by a bicubic upsampling to the target spatial size.
 IV-C Ablation Studies
@@
 IV-C2 Effectiveness of DPA
-To comprehensively evaluate the individual and synergistic effects of components in our proposed Depthwise Position-aware Attention (DPA) module, we perform systematic ablation studies on the Helen test set [1], with detailed results presented in Table II.
+To comprehensively evaluate the individual and synergistic effects of components in our proposed Depthwise Position-aware Attention (DPA) module, we perform systematic ablation studies on the Helen test set, with detailed results presented in Table II.
 Our investigation focuses on three critical elements: the learnable Temperature Generator (T_gen), fixed Temperature Parameter (T_para), and Positional Embedding (P_emb), analyzing their impacts on both reconstruction quality and model complexity.
@@
 The results demonstrate our module’s effectiveness in addressing the unique challenges of facial super-resolution, particularly in preserving identity-critical details while maintaining natural texture synthesis.
+TABLE I: Ablation study of HFR on CelebA test set [3].
+[…]
+w/o HFR 28.16 0.8061
+HFR w/o $\mathit{CS}$ 0.353M 0.802G 28.20 0.8072
+HFR w/ only $\mathrm{DWConv}_{5\times 5}$ 0.316M 0.689G 28.15 0.8058
+HFR w/ only $\mathrm{DWConv}_{7\times 7}$ 0.390M 0.846G 28.18 0.8077
+HFR 0.353M 0.802G 28.21 0.8075
+TABLE II: Ablation study of DPA on Helen test set [41]. The last line denotes the strategy used in our final model.
+[…]
+26.99 0.8034
+0.162M 0.302G 26.97 0.8037
+0.325M 26.93 0.8008
+0.325M 0.302G 27.06 0.8064
+TABLE III: Ablation study of ASSB on the Helen test set [41] by replacing it with three representative alternatives: Self-Attention (SA) [46], Convolutional block (CNN), and Vision State Space Block (VSSB) [8].
+[…]
+w/o ASSB 26.97 0.8039 0.4701 0.2202
+SA 0.184M 0.113G 26.99 0.8035 0.4673 0.2231
+CNN 0.287M 0.387G 27.01 0.8046 0.4691 0.2194
+VSSB 0.269M 0.163G 27.02 0.8042 0.4697 0.2266
+ASSB 0.203M 0.125G 27.06 0.8064 0.4706 0.2167
+TABLE IV: Component-level ablation of ASSB on CelebA test set [3] and Helen test set [41].
+Methods CelebA Helen
+$uparrow PSNR\uparrow / SSIM\uparrow$
+w/o prompt pool 28.16 / 0.8061 26.99 / 0.8035
+w/o routing 28.18 / 0.8069 27.00 / 0.8040
+w/o semantic reorder 28.18 / 0.8074 26.99 / 0.8039
+w/o low-rank prompt 28.19 / 0.8074 27.02 / 0.8045
+Full ASSB 28.21 / 0.8075 27.06 / 0.8064
 IV-C3 Effectiveness of ASSB
-To evaluate the suitability of the Attentive State Space Block (ASSB) for low-frequency modeling in our method, we conduct ablation studies by replacing it with three representative alternatives: self-attention (SA) [47], residual block (CNN) [54], and Vision State Space Block (VSSB) [31].
+To evaluate the suitability of the Attentive State Space Block (ASSB) for low-frequency modeling in our method, we conduct ablation studies by replacing it with three representative alternatives: Self-Attention (SA) [46], Convolutional block (CNN), and Vision State Space Block (VSSB) [8].
 As shown in Table III, our full model with ASSB achieves the best overall performance, indicating superior perceptual quality and structural fidelity.
@@
 In our framework, we integrate ASSB within LFEB to exploit its strength in global structure preservation.
-Unlike Conv, which tends to focus on local or high-frequency features, ASSB’s non-causal state-space formulation excels at capturing smooth, spatially consistent facial regions such as contours and skin tones.
-It’s built-in low-rank prompt decomposition also aids in suppressing high-frequency noise and promotes stable propagation across homogeneous areas.
+Unlike CNN, which tends to focus on local or high-frequency features, ASSB’s non-causal state-space formulation excels at capturing smooth, spatially consistent facial regions such as contours and skin tones.
+Its built-in low-rank prompt decomposition also aids in suppressing high-frequency noise and promotes stable propagation across homogeneous areas.
 Compared to the more general-purpose VSSB, ASSB offers a lighter structure and better frequency selectivity when dealing with low-frequency signals.
@@
 Its integration helps maintain large-scale facial coherence and mitigates structural misalignment—challenges that conventional transformer or Conv-based designs struggle to address.
+We also conduct ablation studies by removing or simplifying individual components in ASSB while keeping the rest of the architecture unchanged.
+These components include the prompt pool, routing strategy, semantic reorder, and low-rank prompt design.
+Together, these ablations cover the key prompt-guided and sequence-organization elements in ASSB, where the semantic reorder variant also reflects the contribution of the non-causal scanning behavior introduced by sequence reorganization.
+As shown in Table IV, simplifying any of these components leads to a performance drop on both datasets.
+This indicates that the improvement brought by ASSB does not arise from naive module stacking, but from the complementary effects of its internal components.
+In particular, the prompt-related design, routing strategy, semantic reorder, and low-rank prompt adaptation all contribute to the final performance.
+Meanwhile, the number of parameters and FLOPs remains nearly unchanged across variants, since these ablations mainly affect internal feature organization and prompt generation strategies rather than the overall backbone width or depth.
+TABLE V: Ablation study of our attention strategy on CelebA test set [3] and Helen test set [41].
+Methods CelebA Helen
+$uparrow PSNR\uparrow / SSIM\uparrow$
+LFEB w/o SEB 28.16 / 0.8067 26.99 / 0.8043
+HFEB w/o DPA 28.04 / 0.8032 26.80 / 0.7981
+w/o offsets 28.17 / 0.8073 27.00 / 0.8050
+Ours 28.21 / 0.8075 27.06 / 0.8064
+TABLE VI: Ablation study of the offset-based alignment mechanism on CelebA test set [3] and Helen test set [41].
+Meehods CelebA Helen
+$uparrow PSNR\uparrow / SSIM\uparrow$
+w/o offset 28.17 / 0.8073 27.00 / 0.8050
+w/ fixed warp 28.20 / 0.8068 27.01 / 0.8047
+w/ zero-offset warp 28.18 / 0.8069 27.00 / 0.8033
+w/ conv alignment 28.18 / 0.8065 27.00 / 0.8049
+Offset (Ours) 28.21 / 0.8075 27.06 / 0.8064
+Figure 5: Comparisons of error maps on our HFR.
+Redder regions indicate larger pixel-wise errors.
+The complete HFR module in our FADPNet leads to the lowest residual errors and shows superior ability to restore high-frequency facial details.
+Figure 6: (a) Ablation studies of frequency-specific modules on the CelebA [3] and Helen datasets [41].
+“swap” is our model in which the low- and high-frequency blocks are swapped, resulting in a noticeable performance drop compared to our original model.
+(b) Feature map visualization and FSR results between our original model and swapped-frequency variant show that our module design for different frequency features produces sharp and accurate facial contours.
 IV-C4 Study of the architecture of low- and high-frequency modules
-To investigate the effectiveness of low- and high-frequency architecture, we conduct an ablation study on the CelebA and Helen datasets by interchanging the positions of the low-frequency enhancement block (LFEB) and high-frequency enhancement block (HFEB).
+To investigate the effectiveness of low- and high-frequency architecture, we conduct an ablation study on the CelebA and Helen datasets by interchanging the positions of LFEB and HFEB.
 As illustrated in Fig.
@@
 Conversely, the swapped version produces noticeably blurrier outputs, especially around facial components like eyes and mouth, further validating the necessity of maintaining the original design of frequency-specific modules.
+TABLE VII: Quantitative comparisons for $\times$8 FSR on CelebA test set [3] and Helen test set [41].
+[…]
+[…]
+Bicubic 23.61 0.6779 0.1821 0.4899 22.95 0.6762 0.1745 0.4912
+FSRNet [13] 27.5M 89.8ms 27.05 0.7714 0.3852 0.2127 25.45 0.7364 0.3482 0.3090
+DIC [14] 22.8M 120.5ms 27.42 0.7840 0.4234 0.2129 26.15 0.7717 0.4085 0.2158
+SPARNet [16] 10.6M 36.6ms 27.73 0.7949 0.4505 0.1995 26.43 0.7839 0.4262 0.2674
+SISN [18] 9.8M 68.0ms 27.91 0.7971 0.4785 0.2005 26.64 0.7908 0.4623 0.2571
+AD-GNN [19] 15.8M 107.9ms 27.82 0.7962 0.4470 0.1937 26.57 0.7786 0.4363 0.2423
+Restormer-M [46] 11.7M 63.2ms 27.94 0.8027 0.4624 0.1933 26.91 0.8013 0.4595 0.2432
+LAAT [20] 22.4M 35.3ms 27.91 0.7794 0.4624 0.1879 26.89 0.8005 0.4569 0.2255
+ELSFace [47] 8.7M 24.0ms 27.96 0.7992 0.4783 0.1927 26.70 0.7910 0.4442 0.2495
+SFMNet [22] 8.6M 49.2ms 27.96 0.7996 0.4644 0.1937 26.86 0.7987 0.4573 0.2322
+SPADNet [48] 13.2M 52.0ms 27.82 0.7966 0.4589 0.1987 26.41 0.7857 0.4295 0.2645
+FreMamba [49] 7.2M 63.0ms 28.13 0.8050 0.4797 0.1807 26.81 0.7967 0.4542 0.2387
+DMNet [34] 8.3M 296.7ms 28.15 0.8045 0.4802 0.1798 26.92 0.7985 0.4664 0.2335
+FreqFormer [33] 9.9M 75ms 28.15 0.8047 0.4883 0.1851 26.82 0.7973 0.4678 0.2290
+FADPNet-S 4.8M 30.0ms 28.13 0.8055 0.4848 0.1807 26.95 0.8015 0.4652 0.2233
+FADPNet 8.6M 39.6ms 28.21 0.8075 0.4876 0.1741 27.06 0.8064 0.4706 0.2167
+Figure 7: Visual comparisons for $\times$8 FSR on CelebA test set [3] and Helen test set [41]. Please zoom in for more details.
+Figure 8: Visual comparisons with existing methods on downstream tasks, like face parsing.
 IV-C5 Study of local-global attention strategy
@@
 Finally, removing the offset mechanism causes a moderate decline in performance, indicating that it contributes to cross-scale feature alignment.
-Figure 10: Visual comparison on SCface [50] of real surveillance scenarios for $\times$8 SR. Our FADPNet can reconstruct clear face components.
-Figure 11: (a) PSNR, FLOPs, and speed tradeoffs on the CelebA test set [4]. (b) PSNR, FLOPs, and Params tradeoffs on the Helen test set [46], showing the balance of our method.
-TABLE VI: Quantitative comparison on face average similarity with existing methods on SCface [50] of real surveillance scenarios.
+To provide a more informative analysis of the offset mechanism, we further compare the proposed offset-based alignment module with several alternatives, including: (1) removing the alignment branch entirely, (2) replacing the learned offsets with a fixed warp, (3) using zero-offset warping, and (4) replacing it with a lightweight convolutional alignment branch.
+As shown in Table VI, replacing the learned offset warp with a fixed warp, zero-offset warp, or convolution-based alignment fails to consistently mitigate cross-scale feature misalignment and leads to inferior overall performance across datasets and metrics.
+This indicates that the improvement does not simply come from adding an extra refinement branch or a generic warping operator, but specifically from the learned offset-based alignment itself.
+Figure 9: Visual comparison on SCface [42] of real surveillance scenarios for $\times$8 SR. Our FADPNet achieves better FSR.
+TABLE VIII: Comparisons on face average similarity with existing methods on SCface [42] of real surveillance scenarios.
 Methods Average Similarity
@@
 SFMNet [22] 0.6240 0.6048 0.6330 0.7158 0.5742
-Ours 0.6782 0.7412 0.7592 0.7714 0.6623
+FADPNet 0.6782 0.7412 0.7592 0.7714 0.6623
+Figure 10: (a) PSNR, FLOPs, and speed tradeoffs on the CelebA test set [3]. (b) PSNR, FLOPs, and Params tradeoffs on the Helen test set [41], showing the balance of our method.
 IV-D Comparison with Other Methods
-In this section, we conduct a comprehensive comparison of our proposed method against state-of-the-art (SOTA) techniques in FSR, encompassing general image SR methods like RCAN [54], Restormer-M [47] and FreqFormer [36], specialized FSR approaches such as FSRNet [13], DIC [14], SPARNet [16], AD-GNN [19], SISN [18], LAAT [20], SFMNet [22], ELSFace [48], and SPADNet [49].
-All methods are trained on the CelebA [4] dataset under identical settings—including preprocessing, loss functions, and optimizer configurations—to eliminate implementation biases and ensure direct attribution of performance differences to architectural innovations.
+We compare our method with existing FSR methods, encompassing general image SR methods like Restormer-M [46], FreqFormer [33] and FreMamba [49], specialized FSR approaches such as FSRNet [13], DIC [14], SPARNet [16], AD-GNN [19], SISN [18], LAAT [20], SFMNet [22], ELSFace [47], SPADNet [48] and DMNet [34].
+All methods are trained on the CelebA [3] dataset under identical settings—including preprocessing, loss functions, and optimizer configurations—to eliminate implementation biases.
 IV-D1 Comparison on the CelebA dataset and the Helen dataset
-We present the quantitative results for the CelebA and Helen test datasets in Table IV.
-The best and second-best results are highlighted in bold and underlined, respectively.
-These results demonstrate the strength of our architecture in balancing structural accuracy and perceptual quality.
-To further evaluate the generalization ability of our method, we directly apply the model trained on CelebA to the Helen test set.
-The quantitative results for $\times$8 FSR are also presented in Table IV, showing that our method comprehensively outperforms other approaches.
-Despite having only 8.6M parameters (similar to SFMNet [22]), ur FADPNet achieves 39.6ms inference speed, which is 1.8× faster than FreqFormer [36] (75ms) and comparable to lightweight methods like SPARNet [16] (36.6ms).
-Compared to ELSFace [48], despite the similar parameter counts (8.6M vs.
-8.7M), FADPNet improves PSNR by 0.25dB on CelebA and 0.36dB on Helen.
-As summarized in Fig.
-7 and Fig.
-8, while competing methods often produce blurred facial contours and distorted features (e.g., ill-defined eyes), our reconstructions successfully preserve sharp edges, facial symmetry, and fine-grained details (e.g., eyeball), closely aligning with the ground-truth HR images.
+Table VII reports quantitative results on CelebA and Helen, with the best and second-best scores marked in bold and underlined, respectively.
+Our model, trained on CelebA and directly applied to Helen, achieves state-of-the-art $\times$8 FSR performance, consistently balancing structural accuracy and perceptual quality.
+With only 8.6M parameters—similar to SFMNet [22]—FADPNet runs at 39.6 ms, 1.8× faster than FreqFormer [33] (75 ms) and comparable to SPARNet [16] (36.6 ms).
+Compared to ELSFace [47], it improves PSNR by 0.25dB on CelebA and 0.36dB on Helen.
+Compared with those recent frequency-aware methods such as FreMamba [49] and DMNet [34], our FADPNet still maintains the best performance and faster inference speed on both CelebA and Helen datasets, further demonstrating the effectiveness of the proposed frequency-dispatched design.
+Our lighter variant, denoted as FADPNet-S (the channel width is reduced, while the overall frequency-aware dual-path framework is preserved), still achieves competitive performance with lower computational cost.
+As also shown in Fig.
+7, competing methods often yield blurred contours or distorted features (e.g., ill-defined eyes), whereas our reconstructions preserve sharp edges, facial symmetry, and fine details (e.g., eyeball), closely aligning with the ground-truth HR images.
+Face parsing results on Helen (as shown in Fig.
+8) further confirm that FADPNet produces the most complete and accurate facial structures.
+Overall, our method outperforms prior methods in both performance (PSNR) and efficiency (Params-FLOPs-Speed) while maintaining robust generalization across datasets.
+IV-D2 Comparison on real-world surveillance
+Restoring facial details from real surveillance imagery is challenging due to LR, uncontrolled lighting, and sensor noise—conditions often underrepresented in synthetic benchmarks.
+We evaluate on SCface [42], which contains naturally degraded surveillance images without manual downsampling.
 As shown in Fig.
-9, we also use the face parsing model [56] to analyze the FSR results of different models on the Helen dataset.
-The visual results clearly demonstrate that our method produces the most complete and accurate facial structure, preserving better face details than competing models.
-In summary, while existing methods struggle with blurred shapes and significant detail loss, our model effectively preserves facial contours and fine details.
-This highlights not only the robustness of our approach across different datasets but also its superiority over SOTA techniques in both performance and efficiency.
-IV-D2 Comparison on real-world surveillance scenarios
-Restoring facial details from real-world surveillance imagery remains a formidable challenge due to inherent low resolution, uncontrolled lighting, and sensor noise—factors that are often inadequately captured in synthetic benchmarks.
-To address this, we conduct experiments using low-quality face images from the SCface dataset [50], which inherently represents real-world surveillance scenarios without manual downsampling.
-As illustrated in Fig.
-10, visual comparisons reveal that existing methods often struggle to reconstruct critical facial features due to the challenges posed by noisy, low-resolution inputs, leading to blurred contours and loss of texture details.
-In contrast, our FADPNet leverages adaptive feature separation and dual frequency enhancement, effectively restoring sharper facial structures, finer textures, and more natural facial symmetry.
-Beyond visual fidelity, we assess the impact on downstream tasks such as face matching, comparing the similarity between restored faces and their corresponding high-definition reference images.
-Table VI presents the quantitative results, demonstrating that our approach consistently achieves higher similarity scores across multiple test cases.
-This confirms its practical applicability in challenging surveillance scenarios.
+9, existing methods frequently produce blurred contours and lose texture details under such conditions.
+In contrast, FADPNet, with adaptive feature separation and dual-frequency enhancement, recovers sharper structures, finer textures, and natural facial symmetry.
+We further assess downstream performance in face matching by comparing restored faces with high-definition references.
+Table VIII shows that FADPNet consistently achieves higher similarity scores, confirming its practical effectiveness in challenging surveillance scenarios.
 IV-D3 Model Complexity Analysis
-Beyond reconstruction quality, we conduct a comprehensive evaluation of model efficiency in terms of parameter count and inference speed, both of which are crucial for devices with limited computing resources.
-At each stage, inputs are dynamically decomposed into high-frequency and low-frequency components, which are then processed by our specialized sub-networks: a CNN branch tailored for capturing high-frequency facial details, and a Mamba-based branch better suited for modeling low-frequency, long-range dependencies.
-This innovative design achieves an optimal balance between computational efficiency and an enhanced representational capacity for high-frequency details.
-By leveraging this structured decomposition, our model adeptly handles complex facial features, leading to improved performance in FSR.
+Beyond FSR quality, we evaluate model efficiency in terms of parameters and inference speed, critical for resource-constrained devices.
+At each stage, inputs are decomposed into high- and low-frequency components, processed respectively by a CNN branch for high-frequency facial details and a Mamba-based branch for low-frequency, long-range dependencies.
+This design balances computational efficiency with strong representation of fine details.
 As shown in Fig.
-11(a) and Fig.
-11(b), our method demonstrates an excellent trade-off between reconstruction accuracy and computational cost.
-Specifically, Fig.
-11(a) illustrates that our approach surpass existing methods over 0.2dB PSNR with lower FLOPs and faster inference speed on the CelebA test set, highlighting its practical efficiency.
-Similarly, Fig.
-11(b) further shows that on the Helen test set, our method maintains superior PSNR while keeping the parameter count relatively low, outperforming most existing methods in both accuracy and model compactness.
-Overall, our approach achieves a well-balanced trade-off between model size, inference speed, and reconstruction quality, making it effective.
+10(a) and Fig.
+10(b), FADPNet achieves superior PSNR with lower FLOPs, faster inference, and fewer parameters compared to existing methods, consistently outperforming them in both accuracy and model cost.
 V Limitations and Future Works